### v1

# Robust Neural Information Retrieval: An Adversarial and Out-of-distribution Perspective

Abstract.
Recent advances in neural information retrieval (IR) models have significantly enhanced their effectiveness over various IR tasks. The robustness of these models, essential for ensuring their reliability in practice, has also garnered significant attention. With a wide array of research on robust IR being proposed, we believe it is the opportune moment to consolidate the current status, glean insights from existing methodologies, and lay the groundwork for future development. We view the robustness of IR to be a multifaceted concept, emphasizing its necessity against adversarial attacks, out-of-distribution (OOD) scenarios and performance variance. With a focus on adversarial and OOD robustness, we dissect robustness solutions for dense retrieval models (DRMs) and neural ranking models (NRMs), respectively, recognizing them as pivotal components of the neural IR pipeline. We provide an in-depth discussion of existing methods, datasets, and evaluation metrics, shedding light on challenges and future directions in the era of large language models. To the best of our knowledge, this is the first comprehensive survey on the robustness of neural IR models, and we will also be giving our first tutorial presentation at SIGIR 2024  [^1] The tutorial slides about robust IR at SIGIR 2024 are available at https://sigir2024-robust-information-retrieval.github.io/.. Along with the organization of existing work, we introduce a Benchmark for robust IR (BestIR), a heterogeneous evaluation benchmark for robust neural information retrieval, which is publicly available at https://github.com/Davion-Liu/BestIR. We hope that this study provides useful clues for future research on the robustness of IR models and helps to develop trustworthy search engines  [^2] A curated list of researches related to robust IR can be found at https://github.com/Davion-Liu/Awesome-Robustness-in-Information-Retrieval..
Robustness, Information Retrieval, Survey
copyright: rightsretainedjournal: TOISjournalvolume: *journalnumber: *article: 1journalyear: 2024publicationmonth: 6ccs: Information systems Retrieval models and ranking
1. Introduction
Information retrieval (IR) plays a pivotal role in numerous real-world applications, such as web search (Ilan, 1998), digital libraries (Chowdhury and Chowdhury, 2003) and e-commerce search (Karmaker Santu et al., 2017). According to the global overview report from Digital 2023, nearly 82% of Internet users between 18 and 64 have used a search engine or web portal in the past month.https://datareportal.com/reports/digital-2023-global-overview-report Specifically, IR is the process of finding and providing relevant information in response to the user query from a large collection of data. Recently, with advances in deep learning, neural IR models have witnessed significant progress (Guo et al., 2020, 2022). With the development of training methodologies such as pre-training (Gao and Callan, 2022; Ma et al., 2022) and fine-tuning (Qu et al., 2021; Zhan et al., 2021; Khattab and Zaharia, 2020), neural IR models have demonstrated remarkable effectiveness in learning query-document relevance patterns.
Why is robustness important in IR?
In real-world deployment of neural IR models, an aspect equally essential as their effectiveness is their robustness. A good IR system must not only exhibit high effectiveness under normal conditions but also demonstrate robustness in the face of abnormal conditions. The natural openness of IR systems makes them vulnerable to intrusion, and the consequences can be severe. For example: (i) Search engines are vulnerable to black hat SEO attacks,https://www.bleepingcomputer.com/news/security/15-000-sites-hacked-for-massive-google-seo-poisoning-campaign/ necessitating significant efforts to curb these infringements.https://www.bbc.com/news/technology-28687513, https://developers.google.com/search/docs/essentials (ii) In addition, search engines are confronted with large amounts of unseen data on a daily basis. The working algorithm needs to be updated frequently to ensure that search effectiveness is maintained.https://developers.google.com/search/news
Figure 1. The statistics of publications related to robust neural information retrieval.
Recently, academics have begun to investigate the robustness of IR systems (Wu et al., 2022b; Thakur et al., 2021; Cohen et al., 2018; Liu et al., 2022a, 2023d; Chen et al., 2023c). As neural networks gain increasing popularity in IR, many studies have found that IR systems that are based on neural networks inherit a wide variety of robustness issues from deep neural network. The field of robust neural IR is garnering increasing attention, as evidenced by the growing number of papers published annually, as depicted in Figure 1. The robustness issues are differently represented in real IR scenarios and raise concerns about deploying neural IR systems into the real world. Therefore, the study of robust neural IR is crucial for building reliable IR systems.
What is the definition of robustness in IR?
In this paper, we begin by defining robustness in the context of IR. In IR, it is well-accepted that, user attention mainly focuses on the Top-$K$ results and increases with higher rankings (Niu et al., 2012). Based on this, we argue that robustness in IR refers to the consistent performance and resilience on Top-$K$ results of an IR system when faced with a variety of unexpected scenarios. Robustness is not a simple concept; it encompasses multiple dimensions, as highlighted by research within the machine learning (ML) community (Shafique et al., 2020; Zhang et al., 2020). Specifically, in the field of IR, we identify several facets of robustness: (i) Performance varianceemphasizes the worst-case performance across different individual queries under the independent and identically distributed (IID) data assumption (Wu et al., 2022b); (ii) Out-of-distribution (OOD) robustnessrefers to the generalizability of an IR model on unseen queries and documents from different distributions of the training dataset (Thakur et al., 2021); and (iii) Adversarial robustnessrefers to the ability of the IR model to defend against malicious adversarial attacks aimed at manipulating rankings (Wu et al., 2022b).
In this survey, our focus is on adversarial robustness and OOD robustness, which have garnered significant attention. (i) For adversarial robustness, studies primarily approach the topic from two angles, i.e., adversarial attack and defense, to enhance the robustness of IR models. (ii) For OOD robustness, the emphasis is on improving the generalizability of IR models to both unseen documents and unseen queries. Additionally, we zero in on two key components of the neural IR framework: first-stage retrieval and the subsequent ranking stage. More specifically, we focus on dense retrieval models (DRMs) and neural ranking models (NRMs) to further explore the aforementioned research perspectives.
The relation to other surveys
There are a number of surveys on robustness in the field of natural language processing (NLP) (Wang et al., 2022c, 2023) and computer vision (CV) (Drenkow et al., 2021; Akhtar et al., 2021). However, the field of IR presents its own unique characteristics: (i) Unlike NLP, which often focuses on individual examples, IR involves ranking a collection of documents, highlighting the need for robustness across the entire list of rankings. (ii) Differing from continuous image data in CV, IR deals with robustness concerning discrete text documents. Consequently, the studies explored in these surveys are not directly transferrable as references within the IR field.
Surveys specific to the IR domain tend to concentrate on areas like pre-training (Zhao et al., 2024; Fan et al., 2022), ranking models (Guo et al., 2020), initial retrieval stages (Guo et al., 2022), and the explainability of IR systems (Anand et al., 2023b). Yet, there’s a noticeable gap in the literature: a dedicated survey that consolidates and introduces research pertaining to robustness in IR is markedly absent.
On top of the work discussed above, this survey adds the following: (i) Our work complements robustness surveys from other fields, taking into account the distinctive characteristics of different domains. (ii) Our survey complements IR surveys by not only prioritizing effectiveness but also emphasizing robustness. Our ultimate aim is to contribute to help advance reliable and efficient IR systems.
Contributions of this survey
We present a compilation of research in IR robustness, summarizing and organizing the content to aid understanding. By categorizing each subfield and delineating its core concepts, we aim to facilitate a deeper understanding of this field and promote the long-term development of robust neural IR. In summary, this paper’s contributions are as follows:
  • A comprehensive overview and categorization: We define robustness in IR by summarizing existing literature and further dividing it into distinct categories.
  • A detailed discussion of methodologies and datasets: We offer a detailed discussion of methodologies, datasets, and evaluation metrics pertinent to each aspect of robustness. Moreover, we integrate the existing datasets mentioned in this survey and propose the BestIR benchmark to facilitate subsequent work.
  • Identification of open issues and future trends: We highlight challenges and potential future trends, particularly in the age of large language models (LLMs).
Organization
Figure 2 depicts the organization of our survey: In Section 2, we introduce the IR task, and give a definition and taxonomy of robustness in IR. In Section 3, we examine adversarial attack and defense tasks, alongside their respective datasets, evaluation criteria, and cutting-edge methodologies. In Section 4, we show two key scenarios for OOD robustness, i.e., OOD generalizability on unseen documents and OOD generalizability to unseen queries, and present specific datasets, evaluation metrics, and methods for solving these scenarios. In Section 5, we describe remaining challenges and emerging opportunities for robustness of IR in the era of LLMs. Finally, in Section 6, we summarize the survey and offer concluding remarks.
Figure 2. Overview of the survey. Section 5 only is partially listed here because of space limitations.
2. Definition and Taxonomy
In this section, we provide a formal definition of robustness in the context of IR and outline the taxonomy pertinent to this domain.
IR task
To provide a clear understanding, we first formalize the IR task. Suppose that $R=\{r_{1},r_{2},\ldots,r_{l}\}$ is the set of relevance levels, where $l$ denotes the number of levels. A total order exists among the relevance labels such that $r_{l}\succ r_{l-1}\succ\cdots\succ r_{1}$, where $\succ$ denotes the order relation. The minimum value of the relevance label is 0, which usually implies no relevance. Suppose that $Q=\{q_{1},q_{2},\ldots,q_{m}\}$ is the set of queries in the training dataset. Each query $q_{i}$ is associated with a list of documents $D_{i}=\{d_{i,1},d_{i,2},\ldots,d_{i,N}\}$ and a list of relevance labels $Y_{i}=\{y_{i,1},y_{i,2},\ldots,y_{i,N}\}$, where $y_{i,j}\in R$ denotes the label of document $d_{i,j}$ and $N$ is the document list size. Then we obtain the training dataset $\mathcal{D}_{\mathrm{train}}=\{(q_{i},D_{i},Y_{i})\}^{m}_{i=1}$.
We use $f$ to denote the IR model, which predicts the relevance score $f(q,d)$ based on a given query $q$ and document $d$. The IR model $f$ is derived by learning from the following objective
(1) $\theta^{*}=\arg\min_{\theta}\mathbb{E}_{(q,D,Y)\sim\mathcal{D}_{\mathrm{train}}}\ \mathcal{L}\left(f(q,d),Y\right),$
where $\theta$ are the parameters of the IR model $f$, and $\mathcal{L}$ is a ranking loss function.
To achieve the IR goal, we employ the ranking function $f$ to generate a score for any query-document pair $(q,d)$, reflecting the relevance degree between $q$ and $d$. This allows us to produce a rank $\pi_{f}\left(q,d\right)$ for each $d\in D$ according to the predicted score. In general, the ranking function $f$ can be further abstracted by the following unified formulation
(2) $f(q,d)=g\left(\psi(q),\phi(d),\eta(q,d)\right),$
where $\psi$ and $\phi$ are representation functions that extract features from $q$ and $d$ respectively, $\eta$ is the interaction function that extracts features from the pair $(q,d)$, and $g$ is an evaluation function that computes the relevance score based on the feature representations. For neural IR models, in most cases, the matching functions $\psi$, $\phi$, $\eta$, and $g$ are encoded in the network structure so that all can be learned from training data (Guo et al., 2020).
The ranking performance of the IR model is usually evaluated by a metric $M$ focus on the top-$K$ ranking results, e.g., Recall@$K$, NDCG@$K$ and MRR@$K$. Given a triple $\left(q,D,Y\right)$ and an IR model $f$, the score of $M$ on query $q$ is calculated by
(3) $M\left(f;\left(q,D,Y\right),K\right)=\sum_{(d,y_{d})\in(D,Y)}y_{d}\cdot h\left(\pi_{f}\left(q,d\right)\right)\cdot\mathbb{I}\left\{\pi_{f}\left(q,d\right)\leq K\right\},$
where $h$ is the mapping function related to ranking, dependent on the specific metric, and $\mathbb{I}\left\{\cdot\right\}$ is an indicator function which is equal to 1 when its condition is satisfied and 0 otherwise.
Further, given a test dataset $\mathcal{D}_{\mathrm{test}}$, the ranking performance $\mathcal{R}_{M}$ against metric $M$ is calculated by
(4) $\mathcal{R}_{M}\left(f;\mathcal{D}_{\mathrm{test}},K\right)=\frac{1}{\left|\mathcal{D}_{\mathrm{test}}\right|}\sum_{\left(q,D,Y\right)\in\mathcal{D}_{\mathrm{test}}}M\left(f;\left(q,D,Y\right),K\right).$
Typically, the ranking performance $\mathcal{R}_{M}$ of the IR model $f$ on the test dataset $\mathcal{D}_{\mathrm{test}}$ against metric $M$ refers to the average evaluation score of $M$.
IR models. In order to balance efficiency and effectiveness, the IR task is usually addressed as a pipelined consisting of a first-stage retrieval stage and re-ranking stage (Chen et al., 2017). (i) The first-stage retrieval identifies a small set of candidate documents from millions of documents. Therefore, in the first-stage retrieval, $D_{i}$ refers to the entire corpus. Considering the efficiency, the neural IR model, represented by the dense IR model (DRM) (Guo et al., 2022), usually adopts a dual-encoder architecture. In this architecture, the interaction function $\eta$ is often null. (ii) The re-ranking stage generates the final ranked list for a query and a small set of candidate documents (Guo et al., 2020), referred to as $D_{i}$ in this stage. To this end, the neural ranking model (NRM) with cross-encoder architecture is often modeled jointly by all the matching functions.
2.1. Definition of Robustness in IR
Robustness refers to the ability to withstand disturbances or external factors that may cause a system to malfunction or provide inaccurate results (Huber, 1981). It is important for practical applications, especially in safety-critical scenarios, e.g., medical retrieval (Averbuch et al., 2004), financial retrieval (Hering, 2017), and private information retrieval (Chor et al., 1998). If, for any reason, the IR system behaves abnormally, the service provider can lose time, manpower, opportunities and even credibility. With the development of deep learning, robustness has received much attention in the fields of computer vision (CV) (Bhojanapalli et al., 2021) and natural language processing (NLP) (Wang et al., 2023). The concerns about model robustness in these fields are mainly focused on the test phase. In this scenario, the model is trained on an unperturbed dataset but tested for its performance when exposed to adversarial examples or OOD data (Goodfellow et al., 2015a; Quionero-Candela et al., 2009).
In IR, the robustness of model in the test phase is also important due to the widespread availability of search engine optimization (SEO) (Gyöngyi and Garcia-Molina, 2005) and the need for models to adapt to unseen data. Hence, in this survey, we follow prior work and only discuss the robustness of a model in the test phase.
Figure 3. The core of robust IR is to protect the stability of the Top-K results.
In IR, it is generally accepted that users focus the vast majority of their attention on the top-$K$ search results, as evidenced by a significant drop in traffic and click-through rates further down the list (Niu et al., 2012; Wu et al., 2022a). Moreover, ranking metrics like MRR (Craswell, 2009) and NDCG (Järvelin and Kekäläinen, 2002), which primarily evaluate the effectiveness of these top-ranking results. The relationship between Top-K result stability and robust IR is shown in Figure 3. Taking SEO as an example, it aims to get its document displayed in the Top-$K$-ranked results. And a robust neural IR model protects its Top-K results from being affected. Consequently, ensuring the integrity and robustness of the top-$K$-ranked results is crucial for deploying IR models into practical web search applications. Based on this, we present a formal definition of top-$K$ robustness in IR by incorporating the original test dataset $\mathcal{D}_{\mathrm{test}}$ from the initial dataset and the unseen test dataset $\mathcal{D}^{*}_{\mathrm{test}}$.
Definition 2.0 (Top-$K$ robustness in information retrieval).
Let $\delta\geq 0$ denote an acceptable error threshold. Given an IR model $f_{\mathcal{D}_{\mathrm{train}}}$ trained on training dataset $\mathcal{D}_{\mathrm{train}}$ with a corresponding testing dataset $\mathcal{D}_{\mathrm{test}}$, an unseen test dataset $\mathcal{D}^{*}_{\mathrm{test}}$, and an acceptable error threshold $\delta$, for the top-$K$ ranking result, if
(5) $\left\lvert\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}};\mathcal{D}_{\mathrm{test}},K\right)-\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}};\mathcal{D}^{*}_{\mathrm{test}},K\right)\right\rvert\leq\delta,$
we consider the model $f_{\mathcal{D}_{\mathrm{train}}}$ to be $\delta$-robust for metric $M$.
Note that in Section 2.2, below $\mathcal{D}^{*}_{\mathrm{test}}$ refers to different test datasets depending on the context: $\mathcal{D}^{\prime}\mathrm{test}$ for test dataset with adversarial examples in adversarial robustness, $\tilde{\mathcal{D}}_{\mathrm{test}}$ for test dataset from new domains in OOD robustness.
Figure 4. A taxonomy of the robustness in IR. In this survey, we pay special attention to adversarial robustness and OOD robustness.
2.2. Taxonomy of Robustness in IR
In IR, robustness threats exist in a wide range of aspects, including adversarial robustness (Wu et al., 2022b), OOD robustness (Thakur et al., 2021) and performance variance (Wu et al., 2022b). The framework of robustness in IR is shown in Figure 4.
2.2.1. Performance variance
Typically, the performance of IR models is first represented by their overall performance on IID data. Recently, it has been recognized that performance stability across IID queries may be compromised when we try to improve the average retrieval effectiveness across all queries (Zhang et al., 2013). Therefore, a robust neural IR model should not only have good retrieval performance on the overall testing queries, but also ensure that the performance on individual queries is not too bad. We give a formal definition of performance variance in IR based on Definition 2.1.
Definition 2.0 (Performance variance of information retrieval).
Given an IR model $f_{\mathcal{D}_{\mathrm{train}}}$ trained on training dataset $\mathcal{D}_{\mathrm{train}}$ with a corresponding IID testing dataset $\mathcal{D}_{\mathrm{test}}$, and an acceptable error threshold $\delta$, for the top-$K$ ranking result, if
(6) $\operatorname{Var}\left(\{M\left(f_{\mathcal{D}_{\mathrm{train}}};\left(q,D,Y\right),K\right)\mid\left(q,D,Y\right)\in\mathcal{D}_{\text{test}}\}\right)\leq\delta,$
where $\mathrm{Var}(\cdot)$ is the variance of the ranking performance of the IR model $f_{\mathcal{D}_{\mathrm{train}}}$ on $\mathcal{D}_{\mathrm{test}}$, then the model $f$ is considered $\delta$-robust in terms of performance variance for metric $M$.
2.2.2. Out-of-distribution robustness
IR models need to cope with a constant stream of unseen data (Wu et al., 2022b). The key behind this challenge is how to adapt the model to new data out of the familiar distribution. There are a variety of OOD scenarios in the field of IR, so the OOD robustness of the model is likewise of wide interest. First, the query entered by the user may be unknown and of varying quality (Penha et al., 2022; Liu et al., 2023b). Then, search engine application scenario migration and incremental new documents will likewise bring in OOD data (Cai et al., 2023; Chen et al., 2023d; Thakur et al., 2021). Manual labeling of unseen data as well as retraining IR models incurs significant resource overhead (Nguyen et al., 2016; Thakur et al., 2021). Therefore, it is crucial how to efficiently train the IR model to have effective performance on unseen data.
In this context, studying the robustness of IR models on OOD data has received a lot of attention. OOD robustness measures the performance of an IR model on unseen queries and documents from different distributions of the training dataset. By introducing the test dataset $\tilde{\mathcal{D}}_{\mathrm{test}}$ in a new domain, we give a formal definition of OOD robustness in IR based on Definition 2.1.
Definition 2.0 (Out-of-distribution robustness of information retrieval).
Given an IR model $f_{\mathcal{D}_{\mathrm{train}}}$, an original dataset with training and test data, $\mathcal{D}_{\mathrm{train}}$ and $\mathcal{D}_{\mathrm{test}}$, drawn from the original distribution $\mathcal{G}$, along with a new test dataset $\tilde{\mathcal{D}}_{\mathrm{test}}$ drawn from the new distribution $\tilde{\mathcal{G}}$, and an acceptable error threshold $\delta$, for the top-$K$ ranking result, if
(7) $\left\lvert\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}};\mathcal{D}_{\mathrm{test}},K\right)-\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}};\tilde{\mathcal{D}}_{\mathrm{test}},K\right)\right\rvert\leq\delta\text{ where}\mathcal{D}_{\mathrm{train}},\mathcal{D}_{\mathrm{test}}\sim\mathcal{G},\tilde{\mathcal{D}}_{\mathrm{test}}\sim\tilde{\mathcal{G}},$
the model $f$ is considered $\delta$-robust against out-of-distribution data for metric $M$.
2.2.3. Adversarial robustness
In a competitive scenario, content providers may aim to promote their products or documents in rankings for specific queries (Kurland and Tennenholtz, 2022). This has provided a market for search engine optimization (SEO) and led to the development of attack techniques against search engines. Traditional attacks on search engines are generally called term spamming. They usually resort to stacking keywords to achieve a ranking boost. In recent years, with the development of deep learning, a number of neural approaches have emerged that attack through more imperceptible perturbations. As search engines evolve, defense methods against term spamming are maturing as well.
Adversarial robustness focuses on the stability of the IR model performance when imperceptible malicious perturbations are added to documents. By introducing a test dataset $\mathcal{D}^{\prime}_{\mathrm{test}}$ with adversarial examples, we give a formal definition of adversarial robustness in IR based on Definition 2.1.
Definition 2.0 (Adversarial robustness in information retrieval).
Given an IR model $f_{\mathcal{D}_{\mathrm{train}}}$ trained on training dataset $\mathcal{D}_{\mathrm{train}}$ with a corresponding testing dataset $\mathcal{D}_{\mathrm{test}}$, a new test dataset $\mathcal{D}^{\prime}_{\mathrm{test}}$ containing adversarial examples, and an acceptable error threshold $\delta$, for the top-$K$ ranking result, if
(8) $\left\lvert\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}};\mathcal{D}_{\mathrm{test}},K\right)-\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}};\mathcal{D}^{\prime}_{\mathrm{test}},K\right)\right\rvert\leq\delta\text{ suchthat }\mathcal{D}^{\prime}_{\mathrm{test}}\leftarrow\mathcal{D}_{\mathrm{test}}\cup D_{\mathrm{adv}},$
where $\mathcal{D}_{\mathrm{test}}\cup D_{\mathrm{adv}}$ denotes injecting the set of all generated adversarial examples $D_{\mathrm{adv}}$ into the original test dataset, and then model $f$ is considered $\delta$-robust against adversarial examples for metric $M$.
In this paper, we pay special attention to two frequently studied types of robustness, i.e., adversarial robustness and OOD robustness. (i) In order to study adversarial robustness, existing work usually proceeds along two lines: adversarial attacks and adversarial defenses. We will discuss these lines in detail in Section 3. (ii) Depending on the types of OOD generalizability in IR models, the existing work can be categorized into the OOD generalizability on unseen documents and the OOD generalizability on unseen queries. We will discuss these directions in detail in Section 4. Research on performance variance mainly includes (Wu et al., 2022b; Zhang et al., 2013, 2014, 2022; Wang, 2009). Relatively little work has been done on this type of robustness.
3. Adversarial Robustness
The deep neural networks have been found vulnerable to adversarial examples that can produce misdirection with human-imperceptible perturbations (Goodfellow et al., 2015b; Ebrahimi et al., 2017). In the field of IR, deep learning-based models are also likely to inherit these adversarial vulnerabilities (Szegedy et al., 2014), which raises widespread concerns about the robustness of neural IR systems. Recall that we give the definition of adversarial robustness in IR elaborated in Section 2.2. Therefore, in this section, we present the specific work of adversarial attack and adversarial defense in adversarial robustness, respectively.
3.1. Overview
In IR, search engine optimization (SEO) has been around since the dawn of the world wide web (Gyöngyi and Garcia-Molina, 2005). This includes white-hat SEO (Goren et al., 2020), which modifies documents in good faith and within the rules and expectations of search engines to optimize the quality of web pages. In contrast, black-hat SEO (Castillo et al., 2011), maliciously exploiting loopholes of search engines, is used to get a site ranking higher in search results. Black-hat SEO creates a poor experience for the audience and is a common concern among site owners.
Traditional web spamming
SEO involves the creation of web pages with no real informational value, designed to manipulate search engine algorithms and distort search results. This deceptive practice lures web users to sites they would not typically visit, undermining the trust relationship between users and search engines and potentially damaging the search engines’ reputation (Baeza-Yates and Ribeiro-Neto, [n. d.]).
web spamming attack. When assessing textual relevance, search engines scrutinize the locations on a web page where query terms appear (Baeza-Yates and Ribeiro-Neto, [n. d.]). These locations are referred to as fields. Common text fields for a page include the document body, the title, the meta tags in the HTML header, and the page’s URL. Additionally, the anchor texts associated with URLs that point to the page are considered part of the page (anchor text field), as they often provide a succinct description of the page’s content. The terms in these text fields play a crucial role in determining the page’s relevance to a specific query, with different weights typically assigned to different fields. Term spamming refers to the manipulation of these text fields to make spam pages appear relevant for certain queries (Castillo et al., 2011).
web spamming detection. web spamming posed a serious threat to search engines in the early days. However, they are also easy to spot and suspect due to their simple implementation (Gyöngyi and Garcia-Molina, 2005). web spamming detection involves identifying and mitigating the practice of term spamming, where keywords are excessively used or manipulated on a web page to artificially inflate its relevance to specific search queries (Robertson and Walker, 1994). Various techniques have been developed to detect the term spamming.
There is a utility-based term spamicity detection method known as OSD (Online Spam Detection) (Zhou et al., 2008; Zhou and Pei, 2009) that has been used to identify adversarial examples. OSD mainly relies on TF-IDF features and has been validated by the Microsoft adCenter (Liu et al., 2022a). It introduces the notion of spamicity to measure how likely a page is spam. Spamicity is a more flexible and user-controllable measure than the traditional supervised classification methods. Using spamicity, online link spam, and term spam can be efficiently detected. However, this detection method can only target changes that add query terms and is prone to misclassification.
Why study adversarial attacks in IR?
With the development of deep learning, neural networks are beginning to be widely used in IR models and have achieved excellent performance. Although traditional web spamming can also have a significant attack effect on neural IR models, this method is difficult to pose a threat due to its ease of detection. However, in the context of black-hat SEO, neural IR models are at risk of being attacked due to the inherent vulnerability inherited from neural networks. Therefore, adversarial attacks are beginning to be studied to expose vulnerability flaws in neural IR models in advance.
Figure 5. Purpose and relationship between adversarial attacks and defenses.
Why study adversarial defense in IR?
To mitigate adversarial attacks, there is a growing body of work that is focusing on adversarial defenses. Adversarial defense focuses on the early hardening of model vulnerabilities discovered by adversarial attacks. Its goal is to obtain robust neural retrieval models to build reliable IR systems.
The relationship between adversarial attacks and defenses is shown in Figure 5. Recently, adversarial robustness has begun to gain widespread attention. Below, we describe these efforts from several perspectives: benchmark datasets, adversarial attacks, and adversarial defense.
Table 1. Benchmark datasets for studying adversarial robustness. #Document denotes the number of documents in corpus; #Qtrain denotes the number of queries for training; #Qdev denotes the number of queries for development; #Qeval denotes the number of queries for evaluation.
Dataset #Document #Qtrain #Qdev #Qeval Reference
MS MARCO document (Nguyen et al., 2016) 3.2M 370K 5,193 5,793 (Wu et al., 2023a; Liu et al., 2022a, 2023c, 2024a; Wu et al., 2022a)
MS MARCO passage (Nguyen et al., 2016) 8.8M 500K 6,980 6,837 (Wang et al., 2022a; Zhong et al., 2023; Wu et al., 2023a; Liu et al., 2023c, 2024b; Wu et al., 2022a; Chen et al., 2023c)
Clueweb09-B (Clarke et al., 2009) 50M 150 - - (Wang et al., 2022a; Liu et al., 2024a)
Natural Questions (Kwiatkowski et al., 2019) 21M 60K 8.8k 3.6k (Liu et al., 2022a; Long et al., 2024; Zhong et al., 2023)
TriviaQA (Joshi et al., 2017) 21M 60K 8.8K 11.3K (Long et al., 2024)
TREC DL19 (Craswell et al., 2020b) - - 43 - (Liu et al., 2022a; Wang et al., 2022a; Parry et al., 2024)
TREC DL20 (Craswell et al., 2021) - - 54 - (Parry et al., 2024)
TREC MB14 (Lin et al., 2013) - - 50 - (Liu et al., 2022a)
ASRC (Raifer et al., 2017) 1,279 - 31 - (Wu et al., 2022b; Goren et al., 2018)
Q-MS MARCO (Liu et al., 2023d) - - 4,000 - (Liu et al., 2023d)
Q-Clueweb09 (Liu et al., 2023d) - - 292 - (Liu et al., 2023d)
DARA (Chen et al., 2023b) 164k 50k 3,490 3,489 (Chen et al., 2023b)
3.2. Benchmark Datasets
In this section, we present the datasets commonly used for studying adversarial robustness. Existing work on attacks and defenses against robustness has focused on experiments on existing IR datasets, as shown in Table 1. All datasets can be found in the BestIR benchmark.
Basic datasets
Some datasets in IR are often adapted for reuse by attack and defense methods as basic datasets. These include MS MARCO document/passage (Nguyen et al., 2016) and Clueweb-09B (Clarke et al., 2009). Some work (Wu et al., 2023a; Liu et al., 2023c, 2024a) performed experiments against attacks and defenses directly on these datasets. For example, prior work usually used the training set of the basic dataset to train an NRM as the target model (Wu et al., 2023a; Liu et al., 2024a). Meanwhile, the queries in the development set are used as target queries, and a portion of the documents in the ranked list of each query are sampled as target documents. Attacks on these target documents measure the performance of adversarial attack methods. We will discuss the specific evaluation methods in Section 3.3.
Expansion of datasets
Some competitions, such as TREC DL19 (Craswell et al., 2020b) and TREC DL20 (Craswell et al., 2021), have provided additional query collections for evaluation against the base dataset. Similarly, these query collections can be used to evaluate the effectiveness of attack methods. Queries in these datasets are often attacked as additional sets of targeted queries. For example, existing work usually trains NRM on the MS MARCO passage dataset and uses the 43 queries in TREC DL19 as the target queries to perform attacks (Chen et al., 2023c; Liu et al., 2022a).
Off-the-shelf datasets
Some research has adapted the above basic datasets to construct new datasets that can be used directly to perform attacks or evaluate defenses. For example, ASRC (Raifer et al., 2017) is based on documents in Clueweb, which are manually modified to generate new adversarial samples for evaluating the model’s adversarial defense abilities. Existing work also used it to study the effects of manual manipulation of documents on the environment IR systems (Vasilisky et al., 2023; Goren et al., 2018). To perform the topic-oriented attack, Liu et al. (2023d) constructed query groups on the same topic based on ORCAS (Craswell et al., 2020a) and the TREC 2012 web Track (Clarke et al., 2009) as a complement to MS MARCO document and Clueweb-09b, respectively. DARA (Chen et al., 2023b) is a dataset for detecting adversarial ranking attacks and includes two types of detection tasks for adversarial documents.
3.3. Adversarial Attack
As neural networks have become increasingly prevalent in IR systems, they have also become a target for adversarial attacks. Studying adversarial attacks can help understand the vulnerability of neural IR models before deploying them in real-world applications, and it can also be used as a surrogate evaluation and support the development of appropriate countermeasures.
0:
A target query $q$, a target document $d$, a query collection $\mathcal{Q}$, a corpus $\mathcal{C}$, a neural IR model
$f$, and a ranking loss function $\tilde{\mathcal{L}}$
0:  An adversarial document $d^{adv}$
1:  if $f$ is a black-box model then
2:     Procedure Surrogate model imitation
3:     Query the model $f$ with $\mathcal{Q}$, and collect ranked lists.
4:     Train the surrogate model $\tilde{f}$ with $\mathcal{Q}$ and the collected ranked lists.
5:     $f=\tilde{f}$
6:  end if
7:  Procedure Adversarial attack
8:  Initialize the adversarial example $d^{adv}$ as a copy of the target document $d$.
9:  for $t\leftarrow 1$ to $\eta$ do
10:     Query the model $f$ with the target query $q$, and collect the ranked list $D$.
11:     Calculate the gradient of $\tilde{\mathcal{L}}$ with respect to the target query $q$ and target document $d$:
12:     $\operatorname{gradient}\leftarrow\nabla_{d}\tilde{\mathcal{L}}\left(f,q,d^{adv},D\right)$
13:     Generate the higher dimensional adversarial perturbation $\rho$:
14:     $\rho\leftarrow\operatorname{normalize}(\operatorname{gradient})$
15:     Mapping high-dimensional perturbations to text space:
16:     $p\leftarrow\rho$
17:     Add textual perturbations:
18:     $d^{adv}\leftarrow d\oplus p$
19:  end for
20:  return  $d^{adv}$
Algorithm 1 Adversarial sample generation for neural IR models
3.3.1. What are the differences between IR attacks and CV/NLP attacks?
Adversarial attacks have undergone significant development in the fields of NLP and CV (Ebrahimi et al., 2017; Goodfellow et al., 2015b), where the most emblematic forms of attack are typically directed at image retrieval and text classification tasks. However, the landscape of adversarial attacks assumes a different contour in the realm of IR: (i) Compared with image retrieval attacks, the IR attacks need to maintain semantic consistency of the perturbed document with the original document by considering the textual semantic similarity, rather than pixel-level perturbations within a fixed range in continuous space; and (ii) Inspired by black-hat SEO, the goal of IR attacks is to inject imperceptible perturbations within a document to improve its ranking for one or multiple specific queries within the entire candidate set or corpus, not inducing classification errors by the model.
Without loss of generality, given a query $q$ and a target document $d$, the goal of generating imperceptible perturbations $p$ to attack against a neural IR model $f$ under top-$K$ ranked results can usually be formalized as:
(9) $\max_{p}\;\left(K-\pi_{f}\left(q,d\oplus p\right)+\lambda\cdot\operatorname{Sim}\left(d,d\oplus p\right)\right),$
where $\pi_{f}\left(q,d\oplus p\right)$ denotes the ranking position of the perturbed document $d\oplus p$ in the ranked list generated by $f$ with respect to query $q$. The $\operatorname{Sim}\left(\cdot\right)$ function measures the similarity between the adversarial example and the original document, both textually as well as semantically. $\lambda$ is a regularization parameter used to balance two goals: keeping the adversarial samples as close as possible to the original document, while allowing the adversarial samples to be ranked as high as possible. Ideally, the adversarial sample $d\oplus p$ should preserve the original semantics of $d$, and be imperceptible to human judges yet misleading to the neural IR models.
3.3.2. What is the attack setting?
Depending on whether the attacker has access to the knowledge of the parameters of the target model, the attack setup can be categorized into two main types (Wu et al., 2023a; Liu et al., 2022a): (i) White-box attacks: Here, the attacker can fully access the model parameters and use the model gradient to directly generate perturbations. (ii) Black-box attacks: Here, the model parameters cannot be obtained; the attacker usually adopts a transfer-based black-box attack paradigm (Cheng et al., 2019): They construct a surrogate white-box model by continuously querying the model and getting the output. Specifically, a surrogate model is trained to simulate the performance of the target model, and then the surrogate model is attacked to the transferability of the adversarial samples to indirectly attack the target model. In IR, attackers can query the target model to obtain a ranked list with rich information. Therefore, the surrogate model often has access to sufficient training data, making this attack effective (Wu et al., 2023a; Liu et al., 2022a). There are various attack methods and target models in IR. We present basic pseudo code to illustrate a fundamental IR attack, in Algorithm 1.
Figure 6. Adversarial retrieval attacks vs. adversarial ranking attacks.
According to the type of target model, attack efforts against IR models can be broadly categorized into two types: adversarial retrieval attacks and adversarial ranking attacks. A schematic diagram of the two types of attack is shown in Figure 6. (i) Adversarial retrieval attacks target the first-stage retrieval, mainly against dense retrieval models; and (ii) Adversarial ranking attacks target the re-ranking stage, mainly against neural ranking models. The methodology for conducting adversarial attacks in IR can differ based on the chosen attack strategy and the target model. A categorization of adversarial attacks in IR is shown in Figure 7.
Adversarial attack Adversarial ranking attack Multi-granular attack (Liu et al., 2024a) Prompt attack (Chen et al., 2023c; Parry et al., 2024) Trigger attack (Song et al., 2020; Goren et al., 2020; Liu et al., 2022a; Song et al., 2022) Word substitution attack (Raval and Verma, 2020; Liang et al., 2018; Wang et al., 2022a; Wu et al., 2023a; Liu et al., 2023d) Adversarial retrieval attack Encoding attack (Boucher et al., 2023) Backdoor attack (Long et al., 2024) Corpus poison attack (Liu et al., 2023c; Zhong et al., 2023; Lin et al., 2023)
Figure 7. Classification of adversarial attacks against neural IR models.
In the following, we introduce evaluation, adversarial retrieval attacks, and adversarial ranking attacks of adversarial attacks in IR.
3.3.3. Evaluation
Evaluation of adversarial attacks includes both attack performance and naturalness performance.
Attack performance
Attack performance mainly refers to the degree of ranking improvement after a target document has been attacked. In general, automatic metrics for attack performance are widely adopted as follows:
  • Attack success rate (ASR/SR) (Wu et al., 2023a; Liu et al., 2022a, 2023d), which evaluates the percentage of target documents successfully boosted under the corresponding target query.
  • Average boosted ranks (Boost/Avg.boost) (Liu et al., 2023c, 2022a, d), which evaluates the average improved rankings for each target document under the corresponding target query.
[…]
  • Normalized ranking shifts rate (NRS) (Wang et al., 2022a; Liu et al., 2023c) evaluates the relative ranking improvement of adversarial examples which are successfully recalled into the initial set with $K$ candidates:
    (10) $\operatorname{NRS}=({\Pi_{d}-\Pi_{d^{adv}}})/{\Pi_{d}}\times 100\%,$
    where $\Pi_{d}$ and $\Pi_{d^{adv}}$ are the rankings of $d$ and $d^{adv}$ respectively, produced by the target IR model.
The effectiveness of an adversary is better with a higher value for all these metrics.
Naturalness performance
Naturalness performance refers primarily to the degree to which a target document is imperceptible to humans after it has been attacked. In general, automatic metrics for naturalness performance and human evaluation are widely adopted as follows:
  • Automatic spamicity detection, which can detect whether target pages are spam or not. The utility-based term spamicity method OSD (Zhou and Pei, 2009) is usually used to detect the adversarial examples.
  • Automatic grammar checkers, i.e., Grammarly [^3] https://app.grammarly.com/ and Chegg Writing, [^4] https://writing.chegg.com/ which calculate the average number of errors in the adversarial examples.
  • Language model perplexity (PPL), which measures the fluency using the average perplexity calculated using a pre-trained GPT-2 model (Radford et al., 2019).
  • Human evaluation, which measures the quality of the attacked documents w.r.t. aspects of imperceptibility, fluency, and semantic similarity (Liu et al., 2023d, 2022a; Wu et al., 2022b).
3.3.4. Adversarial retrieval attack
In this subsection, we introduce the task definition of adversarial retrieval attacks and methods to achieve such attacks.
Task definition
The adversarial retrieval attack task is designed to attack against DRMs. The objective of adversarial retrieval attacks is centered around the manipulation of a document that initially fails to be recalled. By integrating adversarial perturbations, the aim is to ensure that this document is subsequently retrieved by the first-stage neural retrieval model, thereby securing its presence within the candidate set. This approach not only challenges the robustness and reliability of neural retrieval systems but also raises significant questions regarding the integrity of information authenticity (Liu et al., 2023c; Zhong et al., 2023). The goal of adversarial retrieval attacks under top-$K$ ranked results can be formalized as:
(11) $\max_{p}\;\left(K-\operatorname{Recall}_{f}\left(q,d\oplus p\right)+\lambda\cdot\operatorname{Sim}\left(d,d\oplus p\right)\right),$
where $\operatorname{Recall}_{f}\left(q,d\oplus p\right)$ denote the recalled position of the perturbed document $d\oplus p$ generated by the frist stage retrieval model $f$ with respect to query $q$ given the entire corpus. A smaller value of $\operatorname{Recall}$ denotes a higher ranking.
Method
Current retrieval attack methods mainly include corpus poison attacks, backdoor attacks, and encoding attacks.
  • Corpus poison attack. Corpus poisoning attacks construct adversarial samples against a specific query and inject them into the corpus in the inference phase so that they are recalled. MCARA (Liu et al., 2023c) addresses this issue and attempts to mine the vulnerability of the DRM. It introduces the adversarial retrieval attack (AREA) task, which intends to deceive the DRM into retrieving a target document that is outside the initial set of candidate documents. Zhong et al. (2023) adopt the HotFlip method (Ebrahimi et al., 2017) from NLP to iteratively add perturbations in the discrete token space to maximize its similarity to a set of queries. In this way, they maximized the contamination of the corpus by a limited number of documents. MAWSEO (Lin et al., 2023) employs adversarial revisions to achieve real-world cybercriminal objectives, including rank boosting, vandalism detection evasion, topic relevancy, semantic consistency, user awareness (but not alarming) of promotional content, etc.
  • Backdoor attack. Backdoor attacks inject a small proportion of ungrammatical documents into the corpus. When user queries contain grammatical errors, the model will recall the learned triggering pattern and assign high relevance scores to those documents. Long et al. (2024) introduces a novel scenario where the attackers aim to covertly disseminate targeted misinformation, such as hate speech or advertisements, through a retrieval system. To achieve this, they propose a perilous backdoor attack triggered by grammatical errors and ensure that attack models can function normally for standard queries but are manipulated to return passages specified by the attacker when users unintentionally make grammatical mistakes in their queries.
  • Encoding attack. By imperceptibly perturbing documents using uncommon encoded representations, encoding attacks control results across search engines for specific search queries. Boucher et al. (2023) make words look the same as they originally do by adding an offset encoding to them, while the search engines are deceived. The experiment on a mirror of Simple Wikipedia shows that the proposed method can successfully deceive search engines in realistic scenarios.
3.3.5. Adversarial ranking attack
In this subsection, we introduce the task definition of adversarial ranking attacks and existing methods to achieve adversarial ranking attacks.
Task definition
The adversarial ranking attack task is designed to attack against NRMs. The aim of adversarial ranking attacks typically involves introducing adversarial perturbations to a document already present in the candidate set, to manipulate its ranking position either elevating or diminishing it as determined by a neural ranking model. The goal of adversarial ranking attacks under top-$K$ ranked results can usually be formalized as:
(12) $\max_{p}\;\left(K-\operatorname{Recall}_{f}\left(q,d\oplus p\right)+\lambda\cdot\operatorname{Sim}\left(d,d\oplus p\right)\right),$
where $\operatorname{Rank}_{f}\left(q,d\oplus p\right)$ denotes the ranking position of the perturbed document $d\oplus p$ in the final ranked list generated by the NRM $f$ with respect to query $q$. A smaller value of $\operatorname{Rank}$ denotes a higher ranking.
Method. Adversarial ranking attacks against NRMs include word substitution attacks, trigger attacks, and prompt attacks.
  • Word substitution attack. Word substitution attacks typically boost a document’s ranking by replacing a small number of words in the document with synonyms. One common method of white-box word substitution attack is the gradient-based attack, where the attacker uses the gradient of the loss function with respect to the input data to create adversarial examples. These examples are designed to cause the model to make incorrect relevance predictions or rankings. Raval and Verma (2020) present a systematic approach of using adversarial examples to measure the robustness of popular ranking models. They follow a similar approach used in text classification tasks (Liang et al., 2018) and perturb a limited number of tokens (with a minimum of one) in documents, replacing them with semantically similar tokens such that the rank of the document changes. Brittle-BERT (Wang et al., 2022a) adds/replaces a small number of tokens to a highly relevant or non-relevant document to cause a large rank demotion or promotion. Authors find a small set of recurring adversarial words that when added to documents result in successful rank demotion/promotion of any relevant/non-relevant document respectively.
    As for black-box word substitution attacks, in the field of machine learning, Szegedy et al. (2014) find that adversarial examples have the property of cross-model transferability, i.e., the adversarial example generated by a surrogate model can also fool a target model. Black-box attacks in information retrieval usually adopt this transfer-based paradigm due to the excellent performance of imitation of the target model. Wu et al. (2023a) propose the first black-box adversarial attack task against NRM, the word substitution ranking attack (WSRA) task. The WSRA task aims to fool NRMs into promoting a target document in rankings by replacing important words in its text with synonyms in a semantic-preserving way. Based on this task, authors propose a novel pseudo relevance-based adversarial ranking attack method (PRADA), which outperformed web spamming methods by 3.9% in ASR. The WSRA task focuses only on attacks on single query-document pairs and does not take into account the dynamic nature of search engines. Based on this, Liu et al. (2023d) introduce the topic-oriented adversarial ranking attack (TARA) task, which aims to find an imperceptible perturbation that can promote a target document in ranking for a group of queries with the same topic.
  • Trigger attack. Trigger attacks boost document rankings by injecting a generated trigger sentence into a specific location in the document (e.g., the beginning). Song et al. (2020) propose using semantically irrelevant sentences (semantic collisions) as perturbations. They develop gradient-based approaches for generating collisions given white-box access to an NRM. Goren et al. (2020) propose a document manipulation strategy to improve document quality for the purpose of improving document ranking. Liu et al. (2022a) propose a trigger attack method, PAT, empowered by a pairwise objective function, to generate adversarial triggers, which causes premeditated disorderliness with very few tokens. TRAttack (Song et al., 2022) used rewriting existing sentences in the text to improve document ranking with learning ability from the multi-armed bandit mechanism.
  • Prompt attack. Prompt attacks use prompts to guide a language model to generate perturbations based on existing documents to improve document ranking. Chen et al. (2023c) propose a framework called imperceptible document manipulation (IDEM) to produce adversarial documents that are less noticeable to both algorithms and humans. IDEM finds the optimal connect sentence to insert into the document through a language model. Parry et al. (2024) analyze the injection of query-independent prompts, such as “ true” into documents and found that the prompt perturbation method is valid for several sequence-to-sequence relevance models like monoT5 (Nogueira et al., 2020).
  • Multi-granular attack. Multi-granular attacks focus on generating high-quality adversarial examples by incorporating multi-granular perturbations, i.e., word level, phase level, and sentence level. Liu et al. (2024a) propose RL-MARA, a reinforcement learning framework to navigate an appropriate sequential multi-granular ranking attack path. By incorporating word-level, phrase-level, and sentence-level perturbations to generate imperceptible adversarial examples, RL-MARA could increase the flexibility of creating adversarial examples, thereby improving the potential threat of the attack.
3.4. Adversarial Defenses
With the advent of SEO, many defenses were born to counter malicious attacks. In the field of adversarial defenses, much work has been devoted to training robust neural IR models or identifying malicious adversarial examples in advance.
3.4.1. IR defense task
The primary objective of defenses in IR is to maintain, or even enhance, the performance of IR models when the test dataset includes adversarial examples. This involves the implementation of strategies during the training or inference phases: the goal is to ensure that the model’s ability to accurately retrieve relevant documents remains uncompromised, even in the presence of manipulative adversarial perturbations.
Without loss of generality, given a test set $\mathcal{D}_{\mathrm{test}}$ and an adversarial document set $D_{\mathrm{adv}}$, the goal of adversarial defense against an neural IR model $f$ under top-$K$ ranked results can usually be formalized as:
(13) $\max\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}};\mathcal{D}^{\prime}_{\mathrm{test}},K\right)\text{ such that }\mathcal{D}^{\prime}_{\mathrm{test}}\leftarrow\mathcal{D}_{\mathrm{test}}\cup D_{\mathrm{adv}}.$
The adversarial defense task would be in the training or testing phase. In the testing phase, it is usually in the form of attack detection. The training phase is usually in the form of both empirical defense and certified robustness. We present pseudo code to illustrate a fundamental IR defense, as shown in Algorithm 2. The schematic diagram of the three tasks of defense is shown in Figure 8. The specific categorization of adversarial defense in information retrieval is shown in Figure 9.
0:
A query collection $\mathcal{Q}$, a corpus $\mathcal{C}$, a neural IR model $f$, potentially adversarial documents
0:  Safe and robust document rankings
1:  Procedure Attack Detection
2:  Train a detector model $g$ using benign and adversarial examples from $\mathcal{Q}$ and $\mathcal{C}$
3:  for each document $d\in\mathcal{C}$ do
4:     Compute the probability $g(d)$ of $d$ being adversarial
5:     if $g(d)>\text{threshold}$ then
6:        Flag document $d$ as adversarial
7:     end if
8:  end for
9:  Procedure Empirical Defense
10:  Fine-tune $f$ on adversarial examples using augmented training set $\mathcal{Q}^{\prime}$, $\mathcal{C}^{\prime}$
11:  for each training iteration do
12:     Apply random transformations to documents in $\mathcal{C}^{\prime}$ to simulate adversarial perturbations
13:     Update $f$ to minimize loss on transformed documents
14:  end for
15:  Procedure Certified Robustness
16:  Incorporate certified defense methods into $f$ (e.g., randomized smoothing)
17:  for each query $q\in\mathcal{Q}$ do
18:     Compute a robustness certificate for the ranking produced by $f(q)$
19:     if certificate fails then
20:        Adjust $f$ to enhance robustness for $q$
21:     end if
22:  end for
23:  return  Updated and defended neural IR model $f$
Algorithm 2 Adversarial Defense in Neural IR Models
In the following, we introduce the evaluation, attack detection, empirical defense, and certified robustness of adversarial defense in IR.
3.4.2. Evaluation
Adversarial defense assessment includes metrics for the training phase and metrics for the inference phase. Specifically, the defense for the training phase is mainly empirical defense and certified robustness, and the defense for the inference phase is mainly the detection of adversarial samples.
Metrics used in the training phase
The metrics in the training phase are mainly for evaluating the ability of the empirical defense and certified robustness method to maintain the original ranked list in the presence of adversarial samples.
  • CleanMRR@$k$ evaluates Mean Reciprocal Rank (MRR) (Craswell, 2009) performance on a clean dataset.
  • RobustMRR@$k$ (Liu et al., 2024b) evaluates the MRR performance on the attacked test dataset.
  • Attack success rate (ASR) (Liu et al., 2024b) evaluates the percentage of the after-attack documents that are ranked higher than original documents.
  • Location square deviation (LSD) (Liu et al., 2024b) evaluates the consistency between the original and perturbed ranked list for a query, by calculating the average deviation between the document positions in the two lists.
Metrics used in the inference phase
The metrics in the inference phase are mainly used for evaluating the ability of attack detection methods to accurately recognize adversarial samples.
  • Point-wise detection accuracy (Chen et al., 2023b) evaluates the correctness of the detection of whether a single document has been perturbed or not.
  • #DD (Chen et al., 2023b) denotes the average number of discarded documents ranked before the relevant document.
  • #DR (Chen et al., 2023b) denotes the average number of discarded relevant documents.
Figure 8. Overview of attack detection, empirical defense, and certified robustness in neural IR.
Adversarial defense Certified robustness (Wu et al., 2022a) Empirical defense Adversarial training (Wang et al., 2017; Zhang et al., 2021; Zhou et al., 2023; Park and Chang, 2019; Lupart and Clinchant, 2023; Liu et al., 2024b) Data augmentation (Wu et al., 2022a) Attack detection Learning-based detection (Chen et al., 2023b) Language-based detection (Shen et al., 2023; Chen et al., 2023c; Liu et al., 2022a, 2023d) Perplexity-based detection (Song et al., 2020; Chen et al., 2023c; Liu et al., 2022a, 2023c)
Figure 9. Classification of adversarial defenses in neural IR.
3.4.3. Attack detection
While progress in empirical defense and certified robustness aids in training NRMs to be more robust in their defense against potential attacks, the detection of adversarial documents has also been explored.
Perplexity-based detection
Perplexity-based detection mainly uses the difference in the distribution of perplexity between the adversarial samples and the original document under the language model for recognition. Adversarial perturbations applied to original documents can significantly impact the semantic fluency of their content (Liu et al., 2022a, 2023c; Chen et al., 2023c). Song et al. (2020) have developed the perplexity-based detection to counter ranking attacks. Detection involves using a pre-trained language model (PLM) to assess the perplexity of documents, where higher perplexity values indicate less fluent text. Consequently, any document surpassing a certain perplexity threshold is filtered out from consideration.
Language-based detection
Language-based detection primarily uses unnatural language representations to identify adversarial samples. Adversarially generated or modified documents often exhibit grammatical inconsistencies or lack context coherence (Shen et al., 2023; Chen et al., 2023c), since documents that have been maliciously perturbed are usually grammatically incorrect and incoherent. Some work uses grammar checkers, i.e., Grammarly and Chegg Writing, to detect adversarial examples (Liu et al., 2022a, 2023d). The number of grammatical errors in the target document according to the grammar checkers and the quality scoring are used as indicators of linguistic acceptability. Any document deemed to have poor linguistic acceptability is subsequently discarded.
Learning-based detection
The learning-based detection uses neural networks to model the characteristics of adversarial samples, and empirically learns to distinguish the adversarial samples from the clean samples. As mentioned earlier, spamming, perplexity-based, and language-based detectors lack knowledge of the adversarial documents, potentially leading to sub-optimal performance. Consequently, Chen et al. (2023b) introduce two kinds of detection tasks, namely point-wise and list-wise detection, to standardize evaluation processes of the efficacy of adversarial ranking detection methods. They opt to fine-tune the BERT and RoBERTa models using the original and adversarial document pairs present in the training set of the generated dataset. Experimental results demonstrate that a supervised classifier can effectively mitigate known attacks, the detection accuracy can be up to 99.5%, but it performs poorly against unseen attacks.
3.4.4. Empirical defense
Empirical defenses attempt to make models empirically robust to known adversarial attacks; this has been extensively explored in image (Madry et al., 2018) and text classification (Wang et al., 2021). The aim of an empirical defense is to find adversarial examples during training and use them to augment the training set.
Data augmentation
Data augmentation often employs randomized or heuristic methods to transform the training samples, thereby expanding the training data. In the training phase, models that have seen augmented data will have some defense against adversarial samples. There are many data augmentation methods in NLP that achieve good defense performance (Ribeiro et al., 2018; Jia and Liang, 2017).
In IR, Wu et al. (2022a) have been the first to apply data augmentation to NRM defense. They find that data augmentation can reduce the attack success rate to some extent. However, it performs worse than customized defenses for IR. Hence, simply augmenting the training documents (as in NLP) is not a robust defense against attacks in IR.
Adversarial training
Adversarial training is one of the most effective defenses against specific seen attacks. By integrating pre-constructed adversarial samples into training data, the adversarial training has demonstrated a strong defense in both CV and NLP (Madry et al., 2018; Zhu et al., 2019).
In IR, there is a body of work that implements adversarial training by means of adversarial optimization (Wang et al., 2017; Zhang et al., 2021; Zhou et al., 2023), such as GAN (Goodfellow et al., 2020). In this context, the goal is often to simply improve the effectiveness of the IR model. Meanwhile, there is work by Lupart and Clinchant (2023) and Park and Chang (2019) who attempt to use adversarial training for improving robustness. By incorporating adversarial examples into training data, adversarial training has become the de facto defense approach to adversarial attacks against NRMs. However, this defense mechanism is subject to a trade-off between effectiveness and adversarial robustness.
To tackle this issue, Liu et al. (2024b) theoretically define the perturbation invariance of a ranking model and design a perturbation-invariant adversarial training (PIAT) method for ranking models to achieve a better effectiveness-robustness trade-off. Experimental results on several ranking models demonstrate the superiority of PITA compared to earlier adversarial defenses.
3.4.5. Certified robustness
Since empirical defenses are only effective for certain attacks rather than all attacks, competition emerges between adversarial attacks and defense methods. To solve the attack-defense dilemma, researchers resort to certified defenses to make models provably robust to certain kinds of adversarial perturbations. In NLP, Jia et al. (2019) have been the first to propose to certify the robustness of adversarial word substitutions by using Interval Bound Propagation (IBP, Dvijotham et al., 2018).
In the field of IR, Wu et al. (2022a) propose a rigorous and provable certified defense method for NRMs. They define Certified Top-$K$ Robustness for ranking models since users mainly care about the top-ranked results in real-world scenarios. Then, they propose CertDR to achieve certified top-$K$ robustness, based on the idea of randomized smoothing. Their experiments demonstrate that CertDR can significantly outperform state-of-the-art empirical defense methods for ranking models.
4. Out-of-distribution Robustness
In addition to adversarial vulnerability, deep neural networks lack generalizability to OOD data. When faced with data that differs from the distribution of the training data, neural networks struggle to maintain performance. In the field of IR, this problem has begun to be exposed and attract attention as neural IR models are now widely being used (Wu et al., 2022b; Thakur et al., 2021). Existing work focuses on the OOD generalizability of DRMs, since OOD data has a direct impact on the retrieval stage. We refer back to Section 2.2 for a definition of OOD robustness in IR. In this section, we present specific work of OOD generalizability on unseen documents and OOD generalizability on unseen queries, respectively.
Figure 10. OOD generalizability on unseen documents vs. queries in neural IR.
4.1. Overview
IR systems are widely used in fields such as search engines (Ilan, 1998), digital libraries (Chowdhury and Chowdhury, 2003), medical search (Luo et al., 2008), legal search (Maxwell and Schafer, 2008). At the same time, the scenarios that IR systems need to cope with are becoming more complex. For a long time, research work on neural IR models has often been carried out in a narrow independent and identically distributed (IID) setting. In this setting, the model faces homogeneous data during training and testing.
OOD generalizability requirements in IR
With deep neural network models being applied to IR, neural IR models have demonstrated excellent results on many tasks (Guo et al., 2020, 2022). However, in the real world, IR systems need to face more than just a single task or scenario (Thakur et al., 2021). In order to deploy an IR system, it is often necessary to construct a training dataset for the neural IR model in it. This process is usually time-consuming and expensive and hence many IR systems are expected to be able to cope with a wide variety of data not seen during training (Thakur et al., 2021; Liu et al., 2023b; Penha et al., 2022). Therefore, OOD generalizability is a key requirement for contemporary IR systems, given the dynamic nature of user needs and evolving data landscapes. In IR, OOD generalizability is focused on unseen documents and unseen queries as shown in Figure 10.
Why should IR models have the generalizability to unseen documents?
In real-world scenarios, the data landscape is constantly evolving, with new documents and information being generated regularly. It is expensive to annotate each new corpus and retrain the IR models. Therefore, a neural IR model that can generalize well to unseen corpora ensures its relevance and usefulness over time, without requiring constant retraining or fine-tuning. Moreover, in complex real-life scenarios, generalizability to a new corpus helps IR models against distributional shifts or domain-specific biases. This helps to ensure that IR models deliver reliable ranking results irrespective of diverse contexts, qualities, and domains.
Why should IR models have the generalizability to unseen queries?
In practical scenarios, the set of possible queries that users may input is vast and constantly evolving (Penha et al., 2022; Zhuang and Zuccon, 2021). Therefore, IR models must possess the ability to handle queries that were not encountered during training. Without generalizability to unseen queries, IR models risk providing inadequate or irrelevant results, ultimately diminishing user satisfaction and trust in the system (Penha et al., 2022; Campos et al., 2023). The nature of information needs is dynamic and diverse (Guo et al., 2011). Users often express their information needs in varied ways, using different vocabulary, language styles, or even typos (Zhuang et al., 2023; Chen et al., 2022a). A robust neural IR model should be able to understand and accommodate these variations, effectively retrieving relevant information regardless of how the query is formulated. Furthermore, in reality, there is a wide variety of query types that are often not fully or adequately accessible during IR model training (Zhuang and Zuccon, 2022, 2021). But, a good IR system should perform consistently in response to a wide range of query types.
4.2. Benchmark Datasets
In this section, we present the commonly used datasets for studying OOD robustness in IR. All datasets can be found in the BestIR benchmark.
4.2.1. Datasets for unseen documents
The datasets for unseen documents mainly involve adaptation to new corpus datasets and updates to corpus datasets. Existing datasets on unseen documents in IR, as shown in Table 2.
Table 2. Benchmark datasets for unseen documents. #Document denotes the number of documents in corpus; #Qtrain denotes the number of queries for training; #Qdev denotes the number of queries for development; #Qeval denotes the number of queries for evaluation; #Corpus denotes the number of corpora.
Type Dataset #Corpus Publication
Adaptation to new corpus BEIR (Thakur et al., 2021) 18 (Oh et al., 2023; Chen et al., 2023a; Anaya-Isaza and Mera-Jiménez, 2022; Laitz et al., 2024; Anand et al., 2023a; Bonifacio et al., 2022; Li and Gaussier, 2024; Ge et al., 2023; Ram et al., 2022; Dai et al., 2022; Liu et al., 2022b; Reddy et al., 2021; Yu et al., 2022; Izacard et al., 2021; Neelakantan et al., 2022; Liang et al., 2020; Ma et al., 2021b; Chandradevan et al., 2024; Thakur et al., 2021; Ma et al., 2021b; Sachan et al., 2023; Cai et al., 2022; Wang et al., 2022b; Zhan et al., 2022; Xian et al., 2023; Xu et al., 2023; Kang et al., 2024; Xin et al., 2022; Yu et al., 2022; Ni et al., 2022; Lu et al., 2022; Huebscher et al., 2022; Formal et al., 2022; Kasela et al., 2024; Chen et al., 2022c; Lee et al., 2023)
Type Dataset #Document #Qtrain #Qdev #Qeval Publication
CDI-MS (Chen et al., 2023d) 3.2M 370K 5,193 5,793 (Chen et al., 2023d)
Updates to a corpus CDI-NQ (Chen et al., 2023d) 8.8M 500K 6,980 6,837 (Chen et al., 2023d)
LL-LoTTE (Cai et al., 2023) 21M 60K 8.8k 3.6k (Cai et al., 2023)
LL-MultiCPR (Cai et al., 2023) 21M 60K 8.8k 3.6k (Cai et al., 2023)
Adaptation to new corpus datasets
In order to model the migration of models between corpora of different domains, adaptation to new corpus datasets typically aggregates multiple existing domain IR datasets. Of these, BEIR (Thakur et al., 2021) is the most typical, it includes nine different retrieval tasks, such as fact-checking, news retrieval, question answering, entity retrieval, and so on. These 18 datasets are across diverse tasks, diverse domains, task difficulties, and diverse annotation strategies.
Updates to a corpus datasets
In order to follow the updates to a corpus over time, the datasets are mainly constructed by slicing or expanding the existing dataset. For example, to mimic the arrival of new documents in MS MARCO (Nguyen et al., 2016), CDI-MS (Chen et al., 2023d) first randomly sampled 60% documents from the whole corpus as the base documents. Then, it randomly samples 10% documents from the remaining corpus as the new document set, which is repeated 4 times. The web is continuously growing with potential distribution shifts. The updates to a corpus focus on the performance of IR model when updates to a corpus undergoes changes over time.
4.2.2. Datasets for unseen queries
The datasets for unseen queries mainly involve query variation datasets and unseen query type datasets. Existing datasets on unseen queries in IR, are shown in Table 3.
Table 3. Benchmark datasets for unseen queries. #Qeval denotes the number of queries for evaluation.
Type Dataset #Qeval Publication
Query variation DL-Typo (Zhuang and Zuccon, 2022) 60 (Zhuang and Zuccon, 2022)
noisy-MS MARCO (Campos et al., 2023) 5.6k (Campos et al., 2023)
rewrite-MS MARCO (Campos et al., 2023) 5.6k (Campos et al., 2023)
noisy-NQ (Campos et al., 2023) 2k (Campos et al., 2023)
noisy-TQA (Campos et al., 2023) 3k (Campos et al., 2023)
noisy-ORCAS (Campos et al., 2023) 20k (Campos et al., 2023)
variations-ANTIQUE (Penha et al., 2022) 2k (Penha et al., 2022)
variations-TREC19 (Penha et al., 2022) 430 (Penha et al., 2022)
Zhuang and Zuccon (2021) 41k (Zhuang and Zuccon, 2021)
Unseen query type MS MARCO (Nguyen et al., 2016) 15k (Wu et al., 2022b)
L4 (Surdeanu et al., 2008) 10k (Cohen et al., 2018)
Query variation datasets
Query variation datasets contain sets of queries that aim for the same information need but are expressed in various ways, reflecting the natural diversity in how different users might phrase their search queries. Such datasets can include paraphrased queries, queries with typos, order-swapped queries, and queries without stop words. For example, Penha et al. (2022) construct query variation datasets by turning queries from TREC DL19 (Craswell et al., 2020b) and ANTIQUE (Hashemi et al., 2020) into different variants using 4 categories in 10 ways. The importance of query variation datasets lies in their ability to simulate real-world search scenarios, where users often have unique ways of expressing their information needs.
Unseen query type datasets
Unseen query type datasets have queries that are not represented in the training data, either by virtue of their topic or the nature of the information being sought. For example, the MS MARCO dataset (Nguyen et al., 2016) only contains 5 types of queries, i.e., location, numeric, person, description, and entity. The primary purpose of these datasets is to test the generalization ability of IR models to novel, real-world query scenarios that users may present.
4.3. OOD generalizability on unseen documents
As argued above, IR systems need to adapt to different environments and variations in the corpus. However, retraining the neural IR models in each new environment is costly. Previous work has only analyzed the generalizability of existing IR models across different domains (Wu et al., 2022b; Thakur et al., 2021; Ren et al., 2023). In this work, we summarize work on adaptation to a new corpus and updates to a corpus. Figure 11 illustrates how we organize the discussion of different methodologies.
OOD generalizability on unseen documents Updates to a corpus Continual learning for dense retrieval (Cai et al., 2023) Continual learning for generative retrieval (Yoon et al., 2023; Kishore et al., 2023; Chen et al., 2023d; Guo et al., 2024; Mehta et al., 2023a) Adaptation to new corpus Scaling up the model capacity (Ni et al., 2022; Lu et al., 2022) Architectural modifications (Huebscher et al., 2022; Formal et al., 2022; Kasela et al., 2024; Chen et al., 2022c; Lee et al., 2023) Domain modeling (Zhan et al., 2022; Xian et al., 2023; Xu et al., 2023; Kang et al., 2024; Xin et al., 2022; Yu et al., 2022) Data augmentation (Oh et al., 2023; Chen et al., 2023a; Anaya-Isaza and Mera-Jiménez, 2022; Laitz et al., 2024; Anand et al., 2023a; Bonifacio et al., 2022; Li and Gaussier, 2024; Ge et al., 2023; Ram et al., 2022; Dai et al., 2022; Liu et al., 2022b; Reddy et al., 2021; Yu et al., 2022; Izacard et al., 2021; Neelakantan et al., 2022; Liang et al., 2020; Ma et al., 2021b; Chandradevan et al., 2024; Thakur et al., 2021; Ma et al., 2021b; Sachan et al., 2023; Cai et al., 2022; Wang et al., 2022b)
Figure 11. Classification of OOD generalizability on unseen documents.
4.3.1. Definition
Generalizability to unseen documents implies the ability of an IR model to maintain retrieval performance when encountering a new and unfamiliar corpus. In IR, improving the model’s OOD generalizability under unseen documents is mainly reflected in enhancing the retrieval performance of the IR model under various new corpus. Without loss of generality, given a test set $\tilde{\mathcal{D}}_{\mathrm{test}}$ with new corpus $\tilde{\mathcal{C}}$, they draw the new distribution $\tilde{\mathcal{G}}$, the goal of improving the OOD generalizability of a neural IR model $f$ on unseen documents under top-$K$ ranked results can usually be formalized as:
(14) $\max\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}};\tilde{\mathcal{D}}_{\mathrm{test}},K\right)\text{ such that }\mathcal{C}_{\sim\mathcal{G}}\in\mathcal{D}_{\mathrm{train}},\;\tilde{\mathcal{C}}_{\sim\tilde{\mathcal{G}}}\in\tilde{\mathcal{D}}_{\mathrm{test}}.$
Specifically, the new corpus $\tilde{\mathcal{C}}$ may result from two main scenarios with respect to an unseen new corpus and updates to a corpus:
Adaptation to new corpora
Adaptation to new corpora refers to the trained IR models that hard to adapt to a corpus of new domains in the absence of supervised data (Thakur et al., 2021). The goal of improving generalizability on the adaptation to new corpus for a neural IR model $f$ under top-$K$ ranked results can usually be formalized as:
(15) $\max\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}^{o}};\mathcal{D}_{\mathrm{test}}^{n},K\right)\text{ such that }\mathcal{C}_{\sim\mathcal{G}}^{o}\in\mathcal{D}_{\mathrm{train}}^{o},\;\mathcal{C}_{\sim\tilde{\mathcal{G}}}^{n}\in\mathcal{D}_{\mathrm{test}}^{n},$
where $o$ is the original corpus domain and $n$ is the new corpus domain. Among the main solutions for adaptation to a new corpus are data augmentation (Wang et al., 2022b; Izacard et al., 2021; Bonifacio et al., 2022) , domain modeling (Yu et al., 2022), architectural modifications (Kasela et al., 2024), and scaling up the model capacity (Ni et al., 2022).
Updates to a corpus
Updates to a corpus refer to the problem for the trained IR model cannot maintain the ranking performance under continuously arrived new documents (Chen et al., 2023d). The goal of improving generalizability on the updates to a corpus for a neural IR model $f$ under top-$K$ ranked results can be formalized as:
(16) $\max\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}^{\sum^{t}_{0}}};\mathcal{D}_{\mathrm{test}}^{t+1}\right)\text{ such that }\mathcal{C}_{\sim\mathcal{G}}^{\sum^{t}_{0}}\in\mathcal{D}_{\mathrm{train}}^{\sum^{t}_{0}},\;\mathcal{C}_{\sim\tilde{\mathcal{G}}}^{t+1}\in\mathcal{D}_{\mathrm{test}}^{t+1}.$
where $t$ is the time session of each corpus update. The main solution for maintaining the ranking performance is continual learning, but the different paradigms of generative retrieval (GR) (Metzler et al., 2021) and dense retrieval (DR) (Guo et al., 2022) lead to different solutions in the two settings.
In the following, we first introduce the evaluation metrics widely used for OOD generalizability on unseen documents. Then, we detail the existing solutions for adaptation to a new corpus and updates to a corpus, respectively.
4.3.2. Evaluation
OOD generalizability of IR models on unseen documents is mainly measured by the ranking performance under the new corpus. For both adaptation to a new corpus and updates to a corpus, ranking performance is the common evaluation. For updates to a corpus, specifically considering the evaluation of the degree of forgetting of the old corpus.
Metrics for ranking performance
For adaptation to a new corpus and updates to a corpus, the ranking performance of IR models under unseen documents is evaluated by common metrics.
  • NDCG (Järvelin and Kekäläinen, 2002) evaluates the quality of ranked results by measuring the gain of a document based on its position in the ranked list.
  • MRR (Craswell, 2009) evaluates the performance of a ranking result by calculating the average of the reciprocal ranks of the first relevant document answer.
  • HIT (Chen et al., 2023d) evaluates the proportion of times a relevant document is found within a set of top-N ranking results.
  • AP (Mehta et al., 2023b) evaluates the average performance of the ranking performance metrics, overall new domains in adaptation to new corpus, and sessions in updates to a corpus. The ranking performance metric could be any of the above.
Metrics for forgetting degree of the old corpus
updates to a corpus are an ongoing process with many sessions and requires the model to memorize new data without forgetting the old. Therefore, some metrics have been proposed to evaluate the model performance from a time-series perspective.
  • Training Time (Cai et al., 2023) evaluates the total time it takes for the IR model to learn new data while recalling old data.
  • Forgett (Cai et al., 2023) evaluates how much the model forgets at session $t$:
    (17) $\mathrm{Forget_{t}}=\frac{1}{t}\sum_{j=0}^{t-1}\max_{l\in\{0,\ldots,t-1\}}\left(p_{l,j}-p_{t,j}\right),$
    where $p$ is the ranking performance under any common metrics.
  • FWT (Cai et al., 2023) evaluates how well the model transfers knowledge from one session to future sessions:
    (18) $\mathrm{FWT}=\frac{\sum_{i=1}^{j-1}\sum_{j=2}^{T}p_{i,j}}{\frac{T(T-1)}{2}},$
    where $T$ is the total number of sessions.
4.3.3. Adaptation to new corpora
The solution to the adaptation to new corpus challenge involves data augmentation, distributionally robust optimization, and domain-invariant projection.
Data augmentation
Data augmentation involves generating or modifying data in such a way that it bridges the gap between the source domain (the domain where the model was originally trained) and the target domain (the new domain where the model is to be applied). This can include techniques like synthesizing new data examples through transformations that maintain the integrity of the underlying patterns, translating examples from one domain to another, or creating semi-synthetic samples. The GPL (Wang et al., 2022b) uses an unsupervised domain adaptation method Generative Pseudo Labeling, which combines a query generator with pseudo labeling from a cross-encoder. The HyperR (Cai et al., 2022) performs a hyper-prompted training mechanism to enable uniform retrieval across tasks of different domains. There are other works that conduct unsupervised pre-training by using large-scale positive and negative pairs with different data augmentation methods, such as, query generation (Liang et al., 2020; Ma et al., 2021b; Chandradevan et al., 2024; Thakur et al., 2021; Ma et al., 2021b; Sachan et al., 2023), synthetic pre-training (Reddy et al., 2021; Yu et al., 2022; Izacard et al., 2021; Neelakantan et al., 2022), or synthetic relevance labels (Li and Gaussier, 2024; Ge et al., 2023; Ram et al., 2022; Dai et al., 2022; Liu et al., 2022b). Overall, data augmentation can enrich the training set to include more domain-relevant variations, thereby enhancing the model’s ability to generalize across domains.
Very recently, LLMs for data augmentation have significantly enhanced IR models by enabling effective corpus adaptation (Laitz et al., 2024; Anand et al., 2023a; Bonifacio et al., 2022). Anaya-Isaza and Mera-Jiménez (2022) explore various data augmentation strategies combined with transfer learning to improve MRI-based brain tumor detection accuracy. Chen et al. (2023a) develop a cross-domain augmentation network to enhance click-through rate prediction by transferring knowledge between domains with different input features. Oh et al. (2023) propose a prompt-based data augmentation method using generative language models for creating synthetic parallel corpora, improving neural machine translation performance.
Domain modeling
Domain modeling seeks to model the data from both the source and target domains into a common feature space where the differences between the domains are minimized. The idea is to learn a representation of the data that retains the essential information for the task at hand while discarding domain-specific features that might lead to bias or overfitting. By doing so, the model learns to focus on the underlying task without being distracted by differences between the domains. COCO-DR (Yu et al., 2022) use implicit distributionally robust optimization to reweight samples from different source query clusters for improving model robustness over rare queries during fine-tuning. Together with contrastive learning, this approach significantly improves the generalization of DRM over different corpora. There have been many successive efforts to optimize for this problem, including MoDIR (Xin et al., 2022), and ToTER (Kang et al., 2024). Xu et al. (2023) solved the domain OOD challenge by modeling a single passage as multiple units with two objectives. One is the semantic balance between units and the other is the extractability of essential units. Distributionally robust optimization helps in reducing the sensitivity of the model to changes in data distribution, thereby improving its adaptability.
Another way to deal with the new domain data is domain-invariant projection. Zhan et al. (2022) has been the first to use a relevance estimation module for modeling domain-invariant matching patterns and several domain adaption modules for modeling domain-specific features of multiple target corpora. Xian et al. (2023) propose a list-level alignment method, which aligns the distributions of the lists and preserves their list structure. They also demonstrate the superiority of the method on theoretical grounds. The domain-invariant feature space enables the model to perform well on the target domain using knowledge acquired from the source domain, thereby facilitating effective domain adaptation.
Architectural modifications
By optimizing the architecture of an IR model, the model can be made to have good domain adaptability. For instance, hybrid retrieval models have been employed to integrate out-of-domain semantics, enhancing zero-shot capabilities and using core strengths of foundational model features (Chen et al., 2022c; Lee et al., 2023). Additionally, DESIRE-ME (Kasela et al., 2024) uses a mixture-of-experts to tailor retrieval strategies effectively across various domains.
Moreover, methods like employing search agents in hybrid environments or using knowledge distillation with hard negative sampling further support the development of IR systems that maintain high performance in unseen domains (Huebscher et al., 2022; Formal et al., 2022). These strategies collectively enhance the adaptability and effectiveness of retrieval systems across a range of out-of-domain scenarios.
Scaling up the model capacity
Scaling up the model capacity has been identified as a crucial approach that significantly boosts a model’s ability to handle diverse data types and improve retrieval effectiveness across unfamiliar domains.
Ni et al. (2022) explore the impact of enlarging dual encoder architectures. They demonstrate that larger models are not only more capable of handling complex queries but also exhibit enhanced generalization across different domains. In a similar vein, Lu et al. (2022) introduce “Ernie-search” which uses a novel method of self on-the-fly distillation to bridge the gap between cross-encoder and dual-encoder architectures. This technique enhances the dual-encoder’s performance by distilling knowledge from a more powerful cross-encoder, effectively scaling up the retrieval capacity without the direct computational cost typically associated with larger models.
4.3.4. Updates to a corpus
In this scenario, IR models need to be constantly compatible with newly added documents, however, this can lead to catastrophic forgetting problems with old documents. Therefore, continuous learning (De Lange et al., 2021) has become the dominant approach, which aims at adapting the model to the newly added data without losing the ability to understand the old data by quickly adapting to the new unlabeled (little labeled) data.
Continual learning for generative retrieval
In GR, a sequence-to-sequence model is adopted to unify both the indexing and retrieval stages. All document information is encoded into the model parameters. The tight binding of the index to the retrieval module makes updating the index costly. To tackle this challenge, DSI++ (Mehta et al., 2023a) adapts a continual learning method for DSI (Tay et al., 2022) to incrementally index new documents while maintaining the ability to answer user queries related to both previously and newly indexed documents. After that, CorpusBrain++ (Guo et al., 2024) uses the continual learning method on another GR model called CorpusBrain (Chen et al., 2022b). Chen et al. (2023d) propose CLEVER to incrementally index new documents while supporting the ability to query both newly encountered documents and previously learned documents. CLEVER performs incremental product quantization (Jegou et al., 2010) to update a partial quantization codebook, and use a memory-augmented learning mechanism to form meaningful connections between old and new documents. Subsequent work on continuous learning has been devoted to the problem of guaranteeing to updates to a corpus for different GR models, for example, DynamicIR (Yoon et al., 2023) and IncDSI (Kishore et al., 2023).
Continual learning for dense retrieval
In DR, the model needs to learn the representation space of the entire corpus and encode each document into an embedding to serve as the index. Therefore, continual learning for DR should effectively adapt to the evolving distribution with the unlabeled new-coming documents, and avoid re-inferring all embeddings of old documents to efficiently update the index each time the model is updated. L2R (Cai et al., 2023) uses backward-compatible representations to deal with this problem. It first selects diverse support negatives for model training, and then uses a ranking alignment objective to ensure the backward-compatibility of representations.
4.4. OOD generalizability on unseen queries
In the dynamic landscape of information retrieval, where user queries constantly evolve, the ability of IR systems to generalize to unseen queries is paramount. Deep-learning based models, constrained by their training data, often falter when faced with novel query formulations. However, statistics indicate that 15% of daily Google searches are brand new. [^5] https://blog.google/products/search/our-latest-quality-improvements-search/ This challenge becomes particularly pronounced in the context of ever-changing user interests and the introduction of new terminologies. Previous work has analyzed the generalizability of existing IR models under query variants and different query types (Wu et al., 2022b; Penha et al., 2022; Zhuang et al., 2022; Liu et al., 2023b), respectively. In this work, we summarize existing work and how to improve the generalizability of IR models under unseen queries. The specific methodology categorization is shown in Figure 12.
OOD generalizability on unforeseen corpus Unseen query type (Cohen et al., 2018; Bigdeli et al., 2024; Ma et al., 2021a; Lupart et al., 2023; Sciavolino et al., 2021) Query variation Hybird methods (Tasawong et al., 2023; Campos et al., 2023; Sidiropoulos and Kanoulas, 2024; Pan et al., 2023) Contrastive learning methods (Zhuang and Zuccon, 2021; Liu et al., 2023a; Ma et al., 2021a; Sidiropoulos and Kanoulas, 2024) Self-teaching methods (Chen et al., 2022a; Zhuang and Zuccon, 2022; Zhuang et al., 2023)
Figure 12. Classification of OOD generalizability on unseen queries.
4.4.1. Definition
Generalizability to unseen queries indicates the capacity of the IR model to sustain its retrieval performance when confronted with new and unfamiliar query formulations. In the field of IR, enhancing a model’s OOD generalizability with respect to unseen queries mainly involves improving the retrieval accuracy of the IR model across a variety of novel queries. Without loss of generality, given a test set $\tilde{\mathcal{D}}_{\mathrm{test}}$ comprising new queries $\tilde{\mathcal{Q}}$, which introduce a new distribution $\tilde{\mathcal{G}}$, the objective of augmenting the OOD generalizability of a neural IR model $f$ for unseen queries under top-$K$ ranked results can be formalized as:
(19) $\max\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}};\tilde{\mathcal{D}}_{\mathrm{test}},K\right)\text{ such that }\mathcal{Q}_{\sim\mathcal{G}}\in\mathcal{D}_{\mathrm{train}},\tilde{\mathcal{Q}}_{\sim\tilde{\mathcal{G}}}\in\tilde{\mathcal{D}}_{\mathrm{test}}.$
Specifically, the new queries $\tilde{\mathcal{Q}}$ may result from two main scenarios with respect to query variation and unseen query type:
Query variation
Query variations refer to different expressions of the same information need. The way in which information is expressed may impact the effectiveness of IR models (Penha et al., 2022). Some query variants, which introduce additional information, such as query expansion, tend to enhance the retrieval performance (Belkin et al., 1995; Benham et al., 2018). Some introduce noise, such as typos, grammatical errors, and variations in word order, which often challenges the robustness of the IR model (Penha et al., 2022; Campos et al., 2023; Zhuang and Zuccon, 2021). In this paper, we mainly focus on the latter one.
The goal of improving generalizability on the query variation for a neural IR model $f$ under top-$K$ ranked results can usually be formalized as:
(20) $\max\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}};G\left(\mathcal{D}_{\mathrm{test}}\right),K\right)\text{ such that }\mathcal{Q}_{\sim\mathcal{G}}\in\mathcal{D}_{\mathrm{train}},G\left(\mathcal{Q}\right)_{\sim\tilde{\mathcal{G}}}\in G\left(\mathcal{D}_{\mathrm{test}}\right).$
where $G(\cdot)$ is a query variation generator that can generate the query variant based on each query in $\mathcal{Q}$. Among the main solutions for maintaining consistent performance when conformed with query variation are (i) self-teaching method, (ii) contrastive learning method, and (iii) hybrid method.
Unseen query type
Unseen query type refers to unfamiliar query types with new query intents that have not been seen during model training. The main solution for unseen query types is cross-domain regularization. The goal of improving generalizability on an unseen query type for a neural IR model $f$ under top-$K$ ranked results can be formalized as:
(21) $\max\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}^{\tau_{i}}};\mathcal{D}_{\mathrm{test}}^{\tau_{j}},K\right)\text{ such that }\mathcal{Q}_{\sim\mathcal{G}}^{\tau_{i}}\in\mathcal{D}_{\mathrm{train}}^{\tau_{i}},\mathcal{Q}_{\sim\mathcal{G}}^{\tau_{j}}\in\mathcal{D}_{\mathrm{test}}^{\tau_{j}}.$
where $\tau_{i}$ and $\tau_{j}$ are the different types of queries.
In the following, we first introduce evaluation metrics that are widely used for OOD generalizability on unseen queries. Then, we detail existing solutions for query variation and unseen query type, respectively.
4.4.2. Evaluation
OOD generalizability of IR models on unseen queries is mainly measured by the ranking performance under the new query set. In addition to the metrics we mentioned in Section 4.3.2 for measuring ranking performance, there are two other metrics for ranking. There are also specific metrics for unseen query types to evaluate differences in the performance of IR models across query types:
Metrics for ranking performance
In addition to MRR and NDCG, the ranking performance of the IR model under unseen queries is evaluated by other common metrics for query variation and unseen query type:
  • Recall (Chen et al., 2022a) measures the proportion of relevant documents that are successfully retrieved from the total amount of relevant documents available.
  • MAP (Larson, 2010) quantifies the average precision of retrieval across different recall levels, effectively summarizing the precision at each point where a relevant document is retrieved.
Specific metrics for unseen query type
DROOD evaluates the drop rate between the ranking performance on the original type of queries and the ranking performance on the unseen type of queries (Wu et al., 2022b):
(22) $DR_{OOD}=\frac{p_{OOD}-p_{IID}}{p_{IID}},$
where $p_{IID}$ is the ranking performance on original type of queries and $p_{OOD}$ is the ranking performance on unseen type of queries.
4.4.3. Query variation
The solution to the query variation challenge involves (i) self-teaching method, (ii) contrastive learning method, and (iii) hybrid method.
Self-teaching methods
The self-teaching method focuses on distilling the matching capabilities of the IR model on the original clean query to the case of query variants. These methods often align the model output for distillation. Chen et al. (2022a) argue that the drift between query variations and original queries in model representation space affects the subsequent effectiveness of IR models. Based on this, they proposed RoDR, calibrating the in-batch local ranking of query variation to that of the original query for the representation space alignment. Zhuang and Zuccon (2022) also noticed this issue and employed CharacterBERT (El Boukkouri et al., 2020) as the backbone encoder to perform a character-level self-teaching method. This method distilled knowledge from queries without typos into the queries with typos in a character embedding space. ToRoDer (Zhuang et al., 2023) uses a pre-training method that uses bottlenecked information to recover the query variation.
Contrastive learning methods
Contrastive learning methods typically make the model robust to query variants by enhancing the supervision of query variants against the original query as well as positive and negative samples. Zhuang and Zuccon (2021) propose a simple typos-aware training method for BERT-based DRMs and NRMs. During training, this method randomly selected query variants and migrated the supervised signals of the original queries directly for training. By comparing the similarity between a query and its variations and other distinct queries with contrastive learning, Sidiropoulos and Kanoulas (2022) improve the robustness of IR models when encountering typos. After that, MIRS (Liu et al., 2023a) uses a robust contrastive method via injection of [MASK] tokens into query variations and encouraging the representation similarity between the original query and the variation. There is also work on using contrastive learning to enhance the ability of the NRM to handle query variants (Ma et al., 2021a), as well as casting multi-positive contrastive learning (Sidiropoulos and Kanoulas, 2024).
Hybrid methods
Some work enabled IR models to perform better than previously when dealing with query variants through a clever combination of self-teaching and contrastive learning. Tasawong et al. (2023) propose a typo-robust representation learning method that combines contrastive learning with dual self-teaching achieving competitive performance. CAPOT (Campos et al., 2023) introduces a notion of an anchoring loss between the unaltered model and the aligned model and designs a contrastive alignment post-training method to learn a robust model. Sidiropoulos and Kanoulas (2024) argue that previous work did not make full use of positive samples and employed contrastive learning with self-teaching that supports multiple positives. There is also work that uses LLMs to enhance the ability to deal with typos (Pan et al., 2023).
4.4.4. Unseen query type
Usually, part of the unseen query type is caused by unseen documents. Solutions to such problems are presented in Section 14. In this subsection, we present additional solutions for unseen query types.
Wu et al. (2022b) comprehensively analyze the robustness of existing ranking models in the face of unseen query types with 5 types of queries. They found that most NRMs do not generalize well to unseen query types. Even after training on multiple types of queries, NRMs still perform poorly when faced with a new kind of query. Among all ranking models, traditional probabilistic ranking models, such as BM25 (Robertson and Walker, 1994), have the strongest generalizability to OOD query types, while NRMs are the worst. Liu et al. (2023b) analyze the robustness of the generative retrieval model and the dense retrieval model under different query types and found that both models are sensitive to query types.
To improve the generalizability of unseen query types for IR models, Cohen et al. (2018) explore the use of adversarial learning as a regularization technique across different domains within the ranking task framework. By employing an adversarial discriminator and training the NRM across a limited number of domains, the discriminator acts to give negative feedback, thereby preventing the model from adopting domain-specific representations. Bigdeli et al. (2024) propose to integrate two knids of triplet loss functions into neural rankers such that they ensure that each query is moved along the embedding space, through the transformation of its embedding representation, in order to be placed close to its relevant document. In this way, they provided the opportunity to jointly rank documents and difficult queries. There has been some work focusing on the challenges of unseen query types in different scenarios (Ma et al., 2021a; Lupart et al., 2023; Sciavolino et al., 2021).
5. Open Issues and Future Directions
The robust neural IR presents several remaining challenges and opportunities for future research.
5.1. Remaining challenges and issues
In this subsection, we introduce the existing challenges and issues in robust neural IR that need to draw attention to adversarial robustness and OOD robustness, respectively.
5.1.1. Challenges on adversarial robustness in IR
Many other key issues in adversarial robustness in IR have not received much attention.
Penetration attacks against the whole “retrieval-ranking” pipeline
Existing work focuses on attacking the first-stage retrieval (Liu et al., 2023c; Zhong et al., 2023) and re-rank stage (Liu et al., 2022a; Wu et al., 2023a) separately. Penetration attacks focus on exploiting vulnerabilities within the retrieval re-rank pipeline, a critical component in IR systems that ranks results based on relevance to the query. These attacks aim to manipulate rankings by identifying and exploiting weaknesses in the pipeline’s design or its underlying algorithms. This manipulation can result in irrelevant or malicious content being ranked higher than genuine content, compromising the integrity and reliability of the IR system.
Universal attacks
Universal attacks represent a form of adversarial threat that is particularly challenging due to its generalizability across different models and instances (Wallace et al., 2019). Unlike targeted attacks that aim at specific vulnerabilities within a system, universal attacks exploit common weaknesses that are present across a wide range of systems. This makes them incredibly difficult to defend against, as they require solutions that are not just effective for a single model or instance but across the entire spectrum of possible configurations. The development of robust defenses against universal attacks is therefore a significant challenge that demands innovative approaches and a deep understanding of the underlying vulnerabilities.
Dynamic attack scenarios
Most existing work on adversarial attacks is based on static assumptions about search engines (Wu et al., 2023a; Liu et al., 2022a; Chen et al., 2023c). However, search engines operate within a dynamic landscape, which may include changes such as the expansion or reduction of the corpus, and the demotion of documents suspected of spamming in their rankings. In search engines, the search engine results page (SERP) for a query is constantly changing. Existing research indicates that, in the dynamic environment of search engines, current attack methods struggle to maintain a consistent ranking advantage (Liu et al., 2023d). Therefore, designing attack methods that fully consider the dynamic nature of search engines is both practically significant and challenging. At the same time, it’s also worthwhile to explore the development of defense methods that evolve in tandem with updates to the search engine.
Gaming in search engine
The phenomenon of “gaming” in search engines, where individuals or entities manipulate search results for competitive advantage, poses a significant challenge to maintaining the integrity and relevance of search outcomes (Kurland and Tennenholtz, 2022). This competitive manipulation not only undermines the quality of information presented to users but also erodes trust in the search engine’s ability to deliver unbiased and accurate results. As search engines evolve into more sophisticated platforms, so too do the methods employed by those looking to exploit their algorithms for personal gain (Raifer et al., 2017). This ongoing battle between search engines and gamers necessitates the development of more advanced detection and mitigation techniques that can adapt to new gaming strategies, preserving the search engine’s role as a reliable source of information.
Defense against unseen attacks
In existing IR counter defenses, it is often the case that empirical defenses where the attack method is known, can yield good results (Liu et al., 2024b). Whereas in real scenarios, the attack methods are multiple and potentially unknown. Defending against unseen attacks is a paramount challenge in enhancing adversarial robustness in IR systems. These attacks are particularly daunting because they exploit new or unknown vulnerabilities, making traditional defense mechanisms, which are often designed to combat known threats, ineffective. The key to overcoming this challenge lies in the development of adaptive, intelligent systems capable of anticipating potential threats and dynamically adjusting their defense mechanisms in real-time. Achieving this level of adaptability and foresight requires a profound shift in the current paradigms of security in IR, embracing more proactive and predictive approaches.
Defense in practical search engines
Implementing effective defense mechanisms in practical search engines is a balancing act between effectiveness, efficiency, and cost. Effective defense strategies are those that can accurately detect and neutralize threats without significantly impacting the user experience or the relevance of search results. However, the computational resources required for these strategies often come with high costs and can affect the efficiency of the search engine, leading to slower response times and decreased user satisfaction. To address these challenges, search engines are increasingly turning to machine learning and artificial intelligence technologies that can provide scalable and cost-effective solutions. These technologies enable the development of adaptive defense mechanisms that can learn from attack patterns and evolve over time, offering a dynamic approach to security that maintains the delicate balance between protecting the search engine and preserving its performance.
5.1.2. Challenges on OOD robustness in IR
Many key issues in OOD robustness in IR have not received much attention.
Reliance on large-scale data
The reliance on large-scale data sets for training and evaluating IR systems poses significant challenges in ensuring OOD robustness. Large datasets often contain biases and do not adequately represent the diversity of real-world scenarios (Kaplan et al., 2014), leading to models that perform well on seen data but poorly on unseen, OOD examples (Thakur et al., 2021; Yu et al., 2022). Addressing this challenge requires innovative approaches to data collection and model training that prioritize diversity and real-world applicability, ensuring that IR systems remain reliable and effective across a broad range of OOD scenarios.
Lack of harmonized benchmarks for multiple OOD scenarios
A major burden in enhancing OOD robustness in IR is the lack of harmonized benchmarks that accurately reflect the multitude of real-world, OOD scenarios. Without standardized benchmarks, it is difficult to assess the true robustness of IR systems across different contexts and to identify areas for improvement. Developing these benchmarks involves not only capturing a wide range of OOD scenarios but also ensuring that they are representative of the actual challenges faced by IR systems in practice. This effort is crucial for advancing the state of OOD robustness in IR.
Enhancing continuous corpus adaptation for DRMs
Existing work mainly considers the problem of one-shot adaptation of DRMs to new corpora (Thakur et al., 2021; Cohen et al., 2018). However, in real search engines, new documents are constantly added to the search engine and bring in a variety of new domains, which poses a challenge to the continuous adaptation ability of DRMs. To address this, it is essential to develop DRMs that not only quickly adapt to new corpora in a one-shot learning scenario but also continuously learn and adjust as new data is introduced. This requires innovative approaches that can dynamically update the models in an incremental fashion without the need for frequent retraining from scratch. Techniques such as online learning, transfer learning, and meta-learning can play pivotal roles in enhancing the continuous corpus adaptation of DRMs, ensuring they remain effective and relevant in the ever-changing search landscape.
Improving OOD generalization of NRMs
Existing approaches to improve OOD generalization ability mainly target DRMs (Thakur et al., 2021; Yu et al., 2022; Zhan et al., 2022). As an important part of the search engine, the OOD robustness of the NRMs should be optimized along with the DRMs. Enhancing the OOD generalization of NRMs involves developing models that can effectively handle queries that deviate significantly from the training distribution, thereby ensuring the retrieval of relevant and accurate results under a wide range of search scenarios. This challenge requires a multifaceted approach, incorporating advanced machine learning techniques such as robust representation learning, anomaly detection, and domain adaptation strategies. By prioritizing the OOD robustness of NRMs alongside DRMs, search engines can significantly improve their ability to serve high-quality, relevant content to users, even when faced with novel or unexpected queries.
Improving OOD robustness in practical search engines
Enhancing OOD robustness in practical search engines is essential for maintaining their effectiveness, efficiency, and cost-effectiveness. To achieve this, search engines must be able to accurately identify and process queries that fall outside the typical distribution of observed data, ensuring that even uncommon or unseen queries return relevant and useful results. Implementing robust OOD handling mechanisms can significantly improve the quality of search results (Thakur et al., 2021; Liu et al., 2023b), but this often requires sophisticated algorithms that can detect and adapt to OOD queries in real-time (Zhuang and Zuccon, 2022; Penha et al., 2022). While these algorithms can be computationally intensive, leading to higher operational costs, the investment in OOD robustness can ultimately enhance user satisfaction and trust in the search engine (Liu et al., 2023b; Zhuang et al., 2023). Moreover, optimizing these algorithms for efficiency can help mitigate additional costs, ensuring that improvements in OOD robustness also align with the search engine’s operational goals.
5.2. Challenges and Opportunities Posed by Large Language Models
Recently, LLMs have gained attention for their excellent generative abilities. There have been several works applying them to IR with good results. We believe that the introduction of LLMs may bring new robustness problems and also provide new solutions to existing robustness problems.
5.2.1. New challenges to IR robustness from LLMs
Recently, there has been a lot of beginning work on using LLMs for IR tasks. However, these attempts may pose new robustness challenges to LLMs-based IR methods due to the robustness issues of LLMs themselves.
New challenges to adversarial robustness
When applied to IR systems, the adversarial vulnerability of the LLMs themselves is introduced at the same time, as demonstrated by:
  • The vulnerability caused by hallucinations of LLMs. There has been some work using the LLM as a ranking model (Sun et al., 2023). However, LLMs can generate plausible yet factually incorrect or irrelevant information—a phenomenon known as “hallucination”. Existing work has developed attacks against LLM-based ranking models, and vulnerability issues have been found with such models (Liu et al., 2024a). In IR systems, such reliance can lead to the retrieval of misleading or inaccurate data, undermining the trustworthiness and reliability of the system. Addressing these hallucinations is crucial for maintaining the integrity of search results and the credibility of IR systems using LLMs (Shuster et al., 2021; Wu et al., 2023b).
  • Defense costs associated with the scale and opacity of LLMs. LLMs are not only large and complex but also often operate as black boxes with limited transparency into how decisions are made. This opacity complicates efforts to diagnose and mitigate vulnerabilities, making defensive measures both technically challenging and resource-intensive. As LLMs scale, the cost and complexity of implementing robust defensive strategies increase exponentially.
New challenges to OOD robustness
LLMs have shown biases and input sensitivities in existing work (Ji et al., 2023; Ferrara, 2023), and these will affect the OOD robustness of IR systems:
  • Bias in the corpus domain of LLMs. The training process of LLMs may not comprehensively represent the diversity of real-world applications. This training leads to a bias towards the domain characteristics of the training data, which can degrade performance when the model encounters OOD queries or documents in IR tasks. Overcoming this bias is essential for the development of robust, generalizable IR systems with LLMs.
  • Sensitivity of LLMs to query inputs. LLMs can exhibit high sensitivity to slight variations in input (Ji et al., 2023; Ferrara, 2023), potentially leading to significantly different IR outcomes. This sensitivity can be particularly problematic in IR contexts where consistent and predictable retrieval performance is necessary. Enhancing the robustness of LLMs against input variation is a critical area for research and development.
5.2.2. New opportunities for IR robustness via LLMs
While the use of LLMs may introduce new robustness risks, the power of LLMs also provides new ideas for improving robustness. In the field of NLP, many works are used to enhance the robustness of NLP models using LLMs (Moraffah et al., 2024; Helbling et al., 2023; Zeng et al., 2024), but not yet in IR.
New opportunities to adversarial robustness
LLMs hold promise for improving the adversarial robustness of IR systems through their ability to generate and predict adversarial examples:
  • Generating adversarial examples with LLMs. LLMs can be used to create sophisticated adversarial examples to train IR systems, thereby enhancing their ability to withstand malicious attacks. By exposing systems to a wider array of adversarial tactics during training, LLMs can help develop more resilient IR models.
  • IR model defense assisted with LLMs. LLMs can assist in the development of defense mechanisms by predicting and countering adversarial strategies. Using LLMs in simulation environments to anticipate potential attacks allows for the proactive strengthening of IR systems.
New opportunities to OOD robustness
The powerful generation and language understanding ability of LLMs can help to improve the OOD robustness of IR systems:
  • Synthesizing OOD training data with LLMs. LLMs can generate diverse and complex datasets that mirror OOD scenarios (Bonifacio et al., 2022), providing IR systems with the training needed to better handle unfamiliar or novel situations. Such synthetic data can help improve the generalizability and robustness of IR models against OOD inputs.
  • LLMs for OOD detection. Using LLMs’ abilities in language understanding, they can be employed to detect and manage OOD queries effectively. By identifying queries that deviate from the training distribution, LLMs can trigger specialized handling for such cases, thereby enhancing the robustness and reliability of IR systems.
6. Conclusion
The landscape of IR has evolved significantly with the advent of neural methods. This evolution has brought with it new challenges in the robustness of IR systems. These robustness issues undermine the trust of users in search engines, making a critical concern for both researchers and practitioners in the field.
In this survey, we have explored the various forms of robustness in IR, focusing particularly on adversarial robustness and OOD robustness. We have also discussed the remaining challenges of these fields, as well as potential future directions for research. This survey contributes to that journey by providing a structured overview of the current state-of-the-art, offering a roadmap for future research directions, and inspiring continued exploration and innovation in the field. While significant progress has been made in understanding and robustness of IR, there is still much work to be done. As the field continues to evolve, it will be crucial to develop robust defenses against these attacks, to ensure the integrity of search results and maintain the trust of users.
In conclusion, robust neural IR represents a complex and multifaceted problem, but also an opportunity for innovative research and development. As we look forward, the goal of developing IR systems that are not only robust but also adaptable, ethical, and user-centric is essential, promising to redefine the boundaries of what is possible in information retrieval.
Acknowledgements.
This work was funded by the Strategic Priority Research Program of the CAS under Grants No. XDB0680102, the National Key Research and Development Program of China under Grants No. 2023YFA1011602 and 2021QY1701, the National Natural Science Foundation of China (NSFC) under Grants No. 62372431, the Youth Innovation Promotion Association CAS under Grants No. 2021100, the Lenovo-CAS Joint Lab Youth Scientist Project, and the project under Grants No. JCKY2022130C039. This work was also (partially) funded by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl, project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21, which is (partly) financed by the Dutch Research Council (NWO), project ROBUST with project number KICH3.LTP.20.006, which is (partly) financed by the Dutch Research Council (NWO), DPG Media, RTL, and the Dutch Ministry of Economic Affairs and Climate Policy (EZK) under the program LTP KIC 2020-2023, and the FINDHR (Fairness and Intersectional Non-Discrimination in Human Recommendation) project that received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.